### v2

# Generalized discrete Markov spectra

Abstract.
In this paper, we generalize the special subset of the Markov-Lagrange spectrum (and the Markov spectrum) called the discrete Markov spectrum. The discrete Markov spectrum is defined in terms of the Markov numbers, which arise as positive integer solutions to the Markov equation $x^{2}+y^{2}+z^{2}=3xyz.$ Using the tool called snake graphs, originating from cluster algebra theory, we first reconstruct proofs of its properties in a combinatorial framework and then extend it to the generalized setting. We then introduce the generalized discrete Markov spectrum, defined analogously via the generalized Markov numbers, which arise as positive integer solutions to the generalized Markov equation $x^{2}+y^{2}+z^{2}+k_{1}yz+k_{2}zx+k_{3}xy=(3+k_{1}+k_{2}+k_{3})xyz.$ We prove that this generalized spectrum is contained in the Markov-Lagrange spectrum and thus the Markov spectrum.
Key words and phrases:
Markov spectrum, Markov number, Cohn matrix, snake graph
2020 Mathematics Subject Classification:
11J06,11K60,11J70
1. Introduction and Main Results
1.1. Background
The Markov spectrum is the set
$\mathcal{M}:=\left\{\sup_{(x,y)\in\mathbb{Z}^{2}\setminus\{0\}}\frac{\sqrt{D}}{|Q(x,y)|}\;\middle|\;Q(x,y)=ax^{2}+bxy+cy^{2},\;a,b,c\in\mathbb{R},\;D=b^{2}-4ac>0\right\},$
and it is known to have been studied by Markov around 1880. The value
$\mathcal{M}(Q):=\sup_{(x,y)\in\mathbb{Z}^{2}\setminus\{0\}}\frac{\sqrt{D}}{|Q(x,y)|}$
can be understood as a measure of how close the lines defined by the zeros of the quadratic form $Q(x,y)$ comes to lattice points. Markov focused on the part of this spectrum less than 3 and showed that every element in this range can be expressed using a Markov number. More precisely, a Markov number is a positive integer appearing in a solution $(x,y,z)\in\mathbb{Z}_{>0}^{3}$ of the equation
$x^{2}+y^{2}+z^{2}=3xyz,$
and Markov proved the following theorem:
Theorem 1.1 (Markov’s theorem [mar1, mar2]).
Let
$\mathcal{M}_{d}$ $:=\left\{\frac{\sqrt{9n^{2}-4}}{n}\;\middle|\;n\text{ is a Markov number}\right\}$
$\mathcal{M}_{<3}$ $:=\{\mathcal{M}(Q)\mid Q:\text{real indefinite quadratic form and }\mathcal{M}(Q)<3\}\subset\mathcal{M}.$
Then $\mathcal{M}_{d}=\mathcal{M}_{<3}$.
In this paper, we call the set $\mathcal{M}_{d}$ the discrete Markov spectrum. This set is also known to be contained in the Markov-Lagrange spectrum, which is defined as
$\mathcal{L}:=\left\{\limsup_{q\to\infty}\frac{1}{q\|q\alpha\|}\;\middle|\;\alpha\in\mathbb{R}\setminus\mathbb{Q}\right\},$
where $\|q\alpha\|$ denotes the distance from $q\alpha$ to the nearest integer. The quantity
$\mathcal{L}(\alpha):=\limsup_{q\to\infty}\frac{1}{q\|q\alpha\|}$
represents the infimum of those $L$ for which there exist infinitely many rationals $\frac{p}{q}$ satisfying
$\left|\alpha-\frac{p}{q}\right|<\frac{1}{Lq^{2}}.$
This is studied in the context of Diophantine approximation, which concerns approximating irrational numbers by rationals. Hurwitz [hurwitz] pointed out that the subset $\mathcal{L}_{<3}$ of the Markov-Lagrange spectrum below 3 coincides with the discrete Markov spectrum $\mathcal{M}_{d}$, that is,
$\mathcal{L}_{<3}=\mathcal{M}_{<3}=\mathcal{M}_{d}.$
This observation initiated the study of Markov numbers in the context of the Diophantine approximation.
Furthermore, the regions of $\mathcal{L}$ and $\mathcal{M}$ above $3$ have also been studied. These two spectra can be both divided into three regions: the discrete region below $3$, characterized by Markov numbers; the continuous region above the Freiman’s constant $c_{F}$, where
$c_{F}:=\dfrac{2221564096+283748\sqrt{462}}{491993569}\approx 4.5278295661\dots,$
(it is called the Hall’s ray) and the intermediate region between them. We call the continued section $3,c_{F})$ the transition interval. It has been proved that, outside the transition interval, the Markov-Lagrange spectrum and the Markov spectrum coincide completely. It is also known that the intersection of the Markov-Lagrange spectrum with transition interval is strictly contained in the intersection of the Markov spectrum with the same interval, that is,
$\mathcal{L}\cap[3,c_{F})\subsetneq\mathcal{M}\cap[3,c_{F})$
(see [cusick-flahive]). For further details of the history, see [aig, cusick-flahive, reutenauer].
1.2. Main Results
In this paper, we generalize the discrete Markov spectrum in a way corresponding to a certain generalization of Markov numbers. First, consider the $(k_{1},k_{2},k_{3})$-generalized Markov (GM) equation, a generalization of the Markov equation:
$x^{2}+y^{2}+z^{2}+k_{1}yz+k_{2}zx+k_{3}xy=(3+k_{1}+k_{2}+k_{3})xyz,$
where $k_{1},k_{2},k_{3}$ are nonnegative integers. A positive integer appearing in a solution to this equation is called a $(k_{1},k_{2},k_{3})$-generalized Markov number, or $(k_{1},k_{2},k_{3})$-GM number for short. This equation was introduced by [gyo21, bana] for the case $k_{1}=k_{2}=k_{3}=1$ and by [gyomatsu] for the general case in the context of generalized cluster algebras. It has been studied from various perspectives as an equation possessing properties similar to those of the Markov equation (for example, [banaian, banaian-sen, bana-gyo, chen-li, courcy-litman-mizuno, gyoda-maruyama-sato, chen-li2, bao-li, chen-jia]).
One of the main results of this paper is an analogue, in terms of GM numbers, of one direction of the inclusion $\mathcal{M}_{d}\subset\mathcal{M}_{<3}(\subset\mathcal{L})$ stated in Markov’s theorem:
Theorem 1.2 (Theorem 6.1, Corollaries 6.3, 6.6).
We fix $(k_{1},k_{2},k_{3})\in\mathbb{Z}_{\geq 0}^{3}$. Let $n$ be a $(k_{1},k_{2},k_{3})$-GM number, and $i\in\{1,2,3\}$ be the position of $n$ in a positive integer solution to the $(k_{1},k_{2},k_{3})$-GM equation. We set
$\Delta(n,i):=((3+k_{1}+k_{2}+k_{3})n-k_{i})^{2}-4.$
Then, there exist a quadratic irrational $\alpha\in\mathbb{Q}\sqrt{\Delta(n,i)}$ and the corresponding quadratic form $Q=(x-\alpha y)(x-\alpha^{\prime}y)$, where $\alpha^{\prime}$ is the quadratic conjugate of $\alpha$, such that
$\mathcal{L}(\alpha)=\mathcal{M}(Q)=\dfrac{\sqrt{\Delta(n,i)}}{n}.$
In particular, for any $(k_{1},k_{2},k_{3})\in\mathbb{Z}_{\geq 0}^{3}$,
[…]
is contained in $\mathcal{L}$ (thus in $\mathcal{M}$).
Clearly, we have $\mathcal{M}_{d}=\mathcal{M}_{0,0,0}$. Therefore we call $\mathcal{M}_{k_{1},k_{2},k_{3}}$ the $(k_{1},k_{2},k_{3})$-generalized discrete Markov spectrum. In this paper, we compute such $\alpha$ using loops on the once-punctured torus together with their associated periodic continued fractions, and we express the corresponding (generalized) Markov numbers as the numbers of perfect matchings of graphs called snake graphs. A crucial feature of this approach is that it does not rely on Christoffel words (also called Cohn words), which are often used in modern proofs of the Markov theorem, such as those by [aig], [bombieri], and [reutenauer]. In these works, both the Cohn matrices and the strongly admissible sequences, which are indispensable tools to prove Markov’s theorem, are constructed from Christoffel words. This approach simplifies the argument and has thus become the standard method. However, in the generalized setting discussed in this paper, neither the generalized Cohn matrices nor the generalized strongly admissible sequences can be constructed from Christoffel words. Consequently, the proofs that rely on Christoffel words cannot be extended in a straightforward way. In this respect, the snake graphs used in this paper retain more detailed information than Christoffel words and are appropriate to use in the proof of the generalized theorem. We would like to emphasize that, in addition to introducing a generalization of the Markov spectrum, this paper also aims to propose a new proof technique for the classical Markov theorem.
According to Markov’s theorem, the discrete Markov spectrum is characterized by the part of the Markov spectrum below 3. However, it is not known whether the generalized discrete Markov spectrum admits a similar characterization:
Question 1.3 (Question 6.12).
Does there exist a characterization of $\mathcal{M}_{k_{1},k_{2},k_{3}}$ (or the union set of them) that generalizes Markov’s theorem?
We further establish several important properties of the generalized discrete Markov spectra. The first concerns the relationship between the classical discrete Markov spectrum and the $(2,2,2)$-generalized discrete Markov spectrum. It was already known, by [gyomatsu], that classical Markov numbers and $(2,2,2)$-GM numbers are connected through a remarkable relation involving squaring and taking square roots. We now show that a similarly simple and elegant relation holds between their spectra.
Theorem 1.4 (Theorem 6.9).
If $r\in\mathcal{M}_{0,0,0}$, then $3r\in\mathcal{M}_{2,2,2}$. Conversely, if $R\in\mathcal{M}_{2,2,2}$, then $\frac{R}{3}\in\mathcal{M}_{0,0,0}$.
Another important theorem of this paper is the characterization of generalized discrete Markov spectra in the transition interval. The elements of generalized discrete Markov spectra in the transition interval are characterized by the following theorem:
Theorem 1.5 (Theorem 6.10).
Set
$\mathcal{M}^{\prime}:=\bigcup_{(k_{1},k_{2},k_{3})\in\mathbb{Z}^{3}_{\geq 0}}\mathcal{M}_{k_{1},k_{2},k_{3}}.$
Then we have
$\mathcal{M}^{\prime}\cap[3,c_{F})=(\mathcal{M}_{0,0,1}\setminus\{\sqrt{5}\})\cup\{2\sqrt{5}\}.$
The fact that the values (especially those with large numerators) appearing in $(\mathcal{M}_{0,0,1}\setminus{\sqrt{5}})\cup{2\sqrt{5}}$ occur as values of the Markov-Lagrange spectrum and the Markov spectrum is likely a new discovery. However, it is known that there exist elements of the Markov-Lagrange spectrum contained in the transition interval other than these values, i.e.
$\mathcal{M}^{\prime}\cap[3,c_{F})\subsetneq\mathcal{L}\cap[3,c_{F}).$
See Remark 6.11.
Finally, let us turn to the extension of Frobenius’s uniqueness conjecture. The uniqueness conjecture is stated as follows.
Conjecture 1.6 (Conjecture 7.1).
For any Markov number $c$, there exists a unique Markov triple $(a,b,c)$ such that $a\leq b\leq c$.
If we were to extend this conjecture in a straightforward manner to GM numbers, the following question naturally arises:
Question 1.7 (Question 7.3).
For any $(k_{1},k_{2},k_{3})$-GM number $c$, is there a unique $(k_{1},k_{2},k_{3})$-GM triple $(a,b,c)$ such that $a\leq b\leq c$?
However, this question has already been answered in the negative by [gyomatsu]. Therefore, based on the main results, we propose a generalization of the equivalent form of the uniqueness conjecture, for which no counterexamples are currently known (Conjecture 7.6).
1.3. Organization
Section 2 describes properties of the classical discrete Markov spectrum. In this section, we also compare our newly introduced construction of admissible sequences with the conventional construction based on Christoffel words. In Section 3 we introduce the generalized Markov equation and generalized Markov numbers, and in Section 4 we present the generalized Cohn matrix, together with its matrix factorization. Section 5 introduces the notions of generalized Markov length and generalized Markov distance, which play a key role in the proof of Theorem 1.2, and Section 6 provides the proof of Theorem 1.2 and some properties, including Theorems 1.4 and 1.5. In Section 7, we will propose the generalized version of the uniqueness conjecture.
Acknowledgements
The author is deeply grateful to Esther Banaian for communicating Theorem 5.7 and many valuable comments. The author would like to thank the participants of the Markov’s Theorem Seminar at the University of Tsukuba from January 27 to 31, 2025: Kota Saito, Hajime Kaneko, Tadahisa Hamada, Katsuki Ito, Satoru Oshima, and Takafumi Tsurumaki. Special thanks are due to Toshiki Matsusaka, and Shigeki Akiyama for their valuable comments. This work was supported by JSPS KAKENHI Grant Number JP25K17224.
2. Classical Discrete Markov Spectrum
In this section, we summarize various well-known facts about the classical Markov spectrum. They are reorganized so as to remain consistent with the subsequent theory of the generalized discrete Markov spectrum.
2.1. Markov Numbers and Markov Tree
We consider the Markov equation
$x^{2}+y^{2}+z^{2}=3xyz.$
A positive integer solution of the Markov equation is called a Markov triple, and positive integer appearing in a Markov triple is a Markov number. We consider the following binary tree $\mathrm{M}\mathbb{T}$:
  • (1)
    The root vertex is $(1,2,1)$.
  • (2)
    For a vertex $(a,b,c)$, its two children are defined as
    ${(a,b,c)}$${\left(b,\dfrac{b^{2}+c^{2}}{a},c\right).}$${\left(a,\dfrac{a^{2}+b^{2}}{c},b\right)}$
This tree is called the Markov tree. The first few vertices in $\mathrm{M}\mathbb{T}$ are as follows:
${(1,2,1)}$${(2,5,1)}$${(1,5,2)}$${(5,13,1)}$${(2,29,5)}$${(5,29,2)}$${(1,13,5)}$${(13,34,1)\cdots}$${(5,194,13)\cdots}$${(29,433,5)\cdots}$${(2,169,29)\cdots}$${(29,169,2)\cdots}$${(5,433,29)\cdots}$${(13,194,5)\cdots}$${(1,34,13)\cdots}$
It is easy to see that for any $(a,b,c)\in\mathrm{M}\mathbb{T}$, the following holds:
Proposition 2.1 (see [gyoda-maruyama-sato]*Proposition 3.2).
The following statements hold:
  • (1)
    Each vertex $(a,b,c)$ in $\mathrm{M}\mathbb{T}$ is a Markov triple with $b>\max\{a,c\}$.
  • (2)
    Every Markov triple $(a,b,c)$ with $b>\max\{a,c\}$ appears exactly once in $\mathrm{M}\mathbb{T}$.
Moreover, each Markov triple has the following property.
Proposition 2.2 (see [aig]*Corollary 3.4).
For any Markov triple $(a,b,c)$, every pair among $a,b,c$ is relatively prime.
We next introduce the fraction labeling of Markov numbers. To this end, we recall the notions of a Farey triple and the Farey tree. We begin with recalling irreducible fractions in $\mathbb{Q}_{\geq 0}\cup\{\infty\}$.
Definition 2.3.
Let $q\in\mathbb{Q}_{\geq 0}\cup\{\infty\}$ and $n$ and $d\in\mathbb{Z}_{\geq 0}$. The symbol $\frac{n}{d}$ is called the reduced expression of $q$ if $n$ and $d$ are relatively prime and $q=\frac{n}{d}$, where $\frac{n}{d}$ is regarded as $\infty$ when $d=0$ and $n>0$. Moreover, a fraction $\frac{n}{d}$ is said to be irreducible if there exists $q\in\mathbb{Q}_{\geq 0}\cup\{\infty\}$ such that $\frac{n}{d}$ is the reduced expression of $q$.
Definition 2.4.
For $\frac{a}{b}$ and $\frac{c}{d}$, we denote adbcad-bc by $\det\!\left(\frac{a}{b},\frac{c}{d}\right)$. A triple $\left(\frac{a}{b},\frac{c}{d},\frac{e}{f}\right)$ is called a Farey triple if
  • (1)
    $\frac{a}{b},\frac{c}{d},\frac{e}{f}$ are irreducible fractions, and
  • (2)
    $\left|\det\!\left(\frac{a}{b},\frac{c}{d}\right)\right|=\left|\det\!\left(\frac{c}{d},\frac{e}{f}\right)\right|=\left|\det\!\left(\frac{e}{f},\frac{a}{b}\right)\right|=1$.
We define the Farey tree $\mathrm{F}\mathbb{T}$ as follows:
  • (1)
    The root vertex is $\left(\frac{0}{1},\frac{1}{1},\frac{1}{0}\right)$.
  • (2)
    Each vertex $\left(\frac{a}{b},\frac{c}{d},\frac{e}{f}\right)$ has the following two children:
    ${\left(\dfrac{a}{b},\dfrac{c}{d},\dfrac{e}{f}\right)}$${\left(\dfrac{a}{b},\dfrac{a}{b}\oplus\dfrac{c}{d},\dfrac{c}{d}\right)}$${\left(\dfrac{c}{d},\dfrac{c}{d}\oplus\dfrac{e}{f},\dfrac{e}{f}\right).}$
    where $\frac{a}{b}\oplus\frac{c}{d}=\frac{a+c}{b+d}$.
The first few vertices of $\mathrm{F}\mathbb{T}$ are as follows:
${\left(\frac{0}{1},\frac{1}{1},\frac{1}{0}\right)}$${\left(\frac{0}{1},\frac{1}{2},\frac{1}{1}\right)}$${\left(\frac{1}{1},\frac{2}{1},\frac{1}{0}\right)}$${\left(\frac{0}{1},\frac{1}{3},\frac{1}{2}\right)}$${\left(\frac{1}{2},\frac{2}{3},\frac{1}{1}\right)}$${\left(\frac{1}{1},\frac{3}{2},\frac{2}{1}\right)}$${\left(\frac{2}{1},\frac{3}{1},\frac{1}{0}\right)}$${\left(\frac{0}{1},\frac{1}{4},\frac{1}{3}\right)\cdots}$${\left(\frac{1}{3},\frac{2}{5},\frac{1}{2}\right)\cdots}$${\left(\frac{1}{2},\frac{3}{5},\frac{2}{3}\right)\cdots}$${\left(\frac{2}{3},\frac{3}{4},\frac{1}{1}\right)\cdots}$${\left(\frac{1}{1},\frac{4}{3},\frac{3}{2}\right)\cdots}$${\left(\frac{3}{2},\frac{5}{3},\frac{2}{1}\right)\cdots}$${\left(\frac{2}{1},\frac{5}{2},\frac{3}{1}\right)\cdots}$${\left(\frac{3}{1},\frac{4}{1},\frac{1}{0}\right)\cdots}$
Proposition 2.5 (see [aig]*Section 3.2).
The following hold:
  • (1)
    If $\left(r,t,s\right)$ is a Farey triple, then so are $\left(r,r\oplus t,t\right)$ and $\left(t,t\oplus s,s\right)$. In particular, every vertex in $\mathrm{F}\mathbb{T}$ is a Farey triple.
  • (2)
    For every irreducible fraction $t\in(0,\infty)$, there exists a unique Farey triple $F$ in $\mathrm{F}\mathbb{T}$ such that $t$ is the second entry of $F$.
  • (3)
    For $\left(r,t,s\right)$ in $\mathrm{F}\mathbb{T}$, the inequalities $r<t<s$ hold.
We define $n_{t}$ to be the Markov number in $\mathrm{M}\mathbb{T}$ that corresponds to the irreducible fraction $t\in[0,\infty]$ at the corresponding position in $\mathrm{F}\mathbb{T}$.
For any Markov triple $(n_{r},n_{t},n_{s})$ in $\mathrm{M}\mathbb{T}$, we consider an integer $u_{t}$ satisfying the following conditions:
$\begin{cases}n_{r}u_{t}\equiv n_{s}\pmod{n_{t}},\\ 0<u_{t}<n_{t}.\end{cases}$
The integer $u_{t}$ is uniquely determined because $n_{r}$, $n_{t}$, and $n_{s}$ are pairwise relatively prime (Proposition 2.2). Moreover, $u_{t}$ depends only on $t$, since $(r,t,s)$ is uniquely determined by $t$ according to Proposition 2.5 (2). In this way, $u_{t}$ is defined for $t\in(0,\infty)$. We set $u_{\frac{0}{1}}=0$ and $u_{\frac{1}{0}}=1$. We call $u_{t}$ the characteristic number of $t$.
Lemma 2.6.
For any irreducible fraction $t\in[0,\infty]$,
$u_{\frac{1}{t}}=n_{t}-u_{t}.$
Lemma 2.6 is essentially proved in [gyoda-maruyama-sato]*Proposition 7.22 (2) (the case $k=0$). Indeed, if $(r,t,s)\in\mathrm{F}\mathbb{T}$, then $(\frac{1}{s},\frac{1}{t},\frac{1}{r})$ appears in $\mathrm{F}\mathbb{T}$ at the position symmetric to $(r,t,s)$ with respect to the central axis of the tree. Moreover, since $(n_{\frac{1}{s}},n_{\frac{1}{t}},n_{\frac{1}{r}})=(n_{s},n_{t},n_{r})$, the element $v_{t}^{+}$ in [gyoda-maruyama-sato] coincides with $u_{\frac{1}{t}}$.
2.2. Cohn Words, Cohn Matrices, and Cohn Tree
Let $\mathfrak{M}_{3}$ be the free (non-commutative) monoid of rank $3$ generated by $\{p,q,r\}$. We consider the following tree, whose vertices are $\mathfrak{M}_{3}^{3}$:
  • (1)
    The root vertex is $(p,q,r)\in\mathfrak{M}_{3}^{3}$.
  • (2)
    For a vertex $(a,b,c)$, its two children are defined as
    ${(a,b,c)}$${(a,ab,b)}$${(b,bc,c).}$
This tree is called the Cohn word tree, and we denote it by $\mathrm{CoW}\mathbb{T}$. We call the word of an element in $\mathfrak{M}_{3}$ appearing in $\mathrm{CoW}\mathbb{T}$ a Cohn word.
The first few Cohn words in $\mathrm{CoW}\mathbb{T}$ are as follows:
[…]
We now recall the geometric realization of Cohn words. For each irreducible fraction $t\in[0,\infty]$, let $c(t)$ denote the element of $\mathfrak{M}_{3}$ in $\mathrm{CoW}\mathbb{T}$ that corresponds to the position of $t$ in $\mathrm{F}\mathbb{T}$. In what follows, we shall introduce a geometric construction of $c(t)$.
We define a word $c_{t}$ associated with a positive irreducible fraction $t$ as follows. Let $t=\frac{a}{b}\in[0,\infty]$ be an irreducible fraction. Consider the line segment in $\mathbb{R}^{2}$ from $(0,0)$ to $(b,a)$, and denote it by $L_{t}$. We give $L_{t}$ the orientation from $(0,0)$ to $(b,a)$. Each time $L_{t}$ intersects a line of the integer lattice, we take the lattice point lying immediately to the right of the intersection point (with respect to the orientation of $L_{t}$), and record these in order as $p_{1},p_{2},\dots,p_{n}$. Here we set $p_{1}=(0,0)$ and $p_{n}=(b,a)$. If the same lattice point appears more than once, we discard duplicates so that all $p_{i}$ are distinct. Connecting the points $p_{1},p_{2},\dots,p_{n}$ in order by line segments, we assign to each segment a letter according to its slope: - if the slope is 0, we assign $p$; - if the slope is $1$, we assign $q$; - if the slope is $\infty$, we assign $r$.
The sequence of these letters, taken in order, is defined to be $c_{t}$.
Example 2.7.
The word $c_{\frac{2}{5}}$ is $p^{2}qpq$, and $c_{\frac{5}{2}}$ is $qrqr^{2}$. See Figure 1.
ppppqqppqq
qqrrqqrrrr
Figure 1. The word $c_{\frac{2}{5}}$ and $c_{\frac{5}{2}}$
Remark 2.8.
The word $c_{t}$ is also known as a Christoffel word.
We then have the following statement:
Theorem 2.9 (See [aig]*Theorem 7.6).
For any irreducible fraction $t\in[0,\infty]$, $c_{t}=c(t)$ holds.
Let
$P=\begin{bmatrix}3&-1\\ 1&0\end{bmatrix},\quad Q=\begin{bmatrix}5&2\\ 2&1\end{bmatrix},\quad R=\begin{bmatrix}2&1\\ 1&1\end{bmatrix}$.
For any irreducible fraction $t\in[0,\infty]$, we define a matrix $C_{t}$ by
$C_{t}:=c_{t}\big|_{p\mapsto P,\ q\mapsto Q,\ r\mapsto R},$
and call it a Cohn matrix. We define the Cohn tree $\mathrm{Co}\mathbb{T}$ by
$\mathrm{Co}\mathbb{T}:=\mathrm{CoW}\mathbb{T}\big|_{(c_{r},c_{t},c_{s})\mapsto(C_{r},C_{t},C_{s})}.$
Remark 2.10.
The Cohn matrix defined here differs from that in [aig] in the arrangement of its entries. More precisely, the present definition is obtained from the one in [aig] by swapping the diagonal entries.
The first few vertices in $\mathrm{Co}\mathbb{T}$ are as follows:
${\left(\begin{bmatrix}3&-1\\ 1&0\end{bmatrix},\begin{bmatrix}5&{2}\\ 2&1\end{bmatrix},\begin{bmatrix}2&{1}\\ 1&1\end{bmatrix}\right)}$${\left(\begin{bmatrix}3&-1\\ 1&0\end{bmatrix},\begin{bmatrix}13&{5}\\ 5&2\end{bmatrix},\begin{bmatrix}5&{2}\\ 2&1\end{bmatrix}\right)}$${\left(\begin{bmatrix}5&{2}\\ 2&1\end{bmatrix},\begin{bmatrix}12&{7}\\ 5&3\end{bmatrix},\begin{bmatrix}2&{1}\\ 1&1\end{bmatrix}\right)}$${\left(\begin{bmatrix}13&{5}\\ 5&2\end{bmatrix},\begin{bmatrix}75&31\\ 29&12\end{bmatrix},\begin{bmatrix}5&{2}\\ 2&1\end{bmatrix}\right)}$${\left(\begin{bmatrix}3&{-1}\\ 1&0\end{bmatrix},\begin{bmatrix}34&{13}\\ 13&5\end{bmatrix},\begin{bmatrix}13&{5}\\ 5&2\end{bmatrix}\right)}$${\left(\begin{bmatrix}12&{7}\\ 5&3\end{bmatrix},\begin{bmatrix}31&{19}\\ 13&8\end{bmatrix},\begin{bmatrix}2&{1}\\ 1&1\end{bmatrix}\right)}$${\left(\begin{bmatrix}5&{2}\\ 2&1\end{bmatrix},\begin{bmatrix}70&{41}\\ 29&17\end{bmatrix},\begin{bmatrix}12&{7}\\ 5&3\end{bmatrix}\right)}$
Theorem 2.11.
For any irreducible fraction $t\in[0,\infty]$, we have
$C_{t}=\begin{bmatrix}3n_{t}-u_{t}&\dfrac{3u_{t}n_{t}-u_{t}^{2}-1}{n_{t}}\\[6.0pt] n_{t}&u_{t}\end{bmatrix}$.
In particular, we have $\det(C_{t})=1$ and $\mathrm{tr}(C_{t})=3n_{t}$.
A proof of Theorem 2.11 can be found in [aig]*Theorem 4.13. In [aig], only the case $t\in[0,1]$ is treated, but the same argument applies to $t\in(1,\infty]$ as well.
2.3. Continued Fraction Matrix Decompositions of Cohn Matrices
In this subsection, we decompose a Cohn matrix into a product of fundamental matrices using geometric information. We consider the square lattice in $\mathbb{R}^{2}$ (including its grid lines), subdivided by lines of slope 1-1 that pass through lattice points. We denote this modified lattice by $\widetilde{\mathbb{R}^{2}}$.
For a given irreducible fraction $t=\frac{a}{b}\in(0,\infty)$, consider again the line segment in $\widetilde{\mathbb{R}^{2}}$ from $(0,0)$ to $(b,a)$, denoted $L_{t}$.
Figure 2. The line segment $L_{t}$ for $t=\frac{2}{5}$
We now define $\overline{L_{t}}$ to be the line segment obtained by shifting $L_{t}$ slightly to the left. We regard that the left endpoint passes through the lower-left edge, whereas the right endpoint remains within the upper-right edge (although the subtle difference in how we treat the two endpoints is not very important in this section, it becomes significant in Section 4 when we assign signs to edges in $\widetilde{\mathbb{R}^{2}}$). In Figure 3, the situation described above is illustrated in a deformed manner for better visual clarity. The same type of deformation will be used in all subsequent figures.
Figure 3. The line segment $\overline{L_{t}}$ associated with $t=\frac{2}{5}$
Remark 2.12.
We can also obtain $\overline{L_{t}}$ by using an one-punctured torus: project $L_{t}$ onto an arc on the triangulated once-punctured torus, slightly push the arc upward to obtain a loop, and view $\overline{L_{t}}$ as the embedding of that loop back into $\widetilde{\mathbb{R}^{2}}$. Figure 4 shows the case where $t=\frac{1}{1}$.
[…]
Figure 4. Interpretation $\overline{L_{t}}$ as loop on torus for $t=\frac{1}{1}$
Using $\overline{L_{t}}$, we construct an integer sequence for any irreducible fraction $t\in(0,\infty)$. To this end, we will introduce the sign rule of triangles in $\widetilde{\mathbb{R}^{2}}$ associated a curve segment.
Let $\gamma$ be a curve segment on $\widetilde{\mathbb{R}^{2}}$. In this paper, a curve segment $\gamma$ is assumed to satisfy the following conditions:
  • the interior of $\gamma$ does not intersect any lattice point of $\widetilde{\mathbb{R}^{2}}$;
  • whenever $\gamma$ meets an edge of $\widetilde{\mathbb{R}^{2}}$, it intersects the edge transversely;
  • $\gamma$ has only finitely many self-intersections.
Definition 2.13.
Let $\gamma$ be an oriented curve segment on $\widetilde{\mathbb{R}^{2}}$. For each triangle in $\widetilde{\mathbb{R}^{2}}$ that intersects $\gamma$, assign a sign in $\{+,-\}$ as follows:
  • (i)
    Assign a sign ()(-) to each triangle whose left-hand piece, obtained by cutting it with $\gamma$, is a quadrilateral (see Figure 5):
    $\mathbf{<}$
    $\mathbf{<}$
    $\mathbf{>}$
    $\mathbf{>}$
    $\mathbf{>}$
    $\mathbf{<}$
    Figure 5. Right-angled triangles with -
  • (ii)
    Assign a sign $(+)$ to all other triangles (see Figure 6):
    $\mathbf{>}$
    $\mathbf{>}$
    $\mathbf{<}$
    $\mathbf{<}$
    $\mathbf{<}$
    $\mathbf{>}$
    Figure 6. Right-angled triangles with ++
This rule is called the triangle-crossing rule of $\gamma$.
Next, we define the strongly admissible sequence s(t)s(t) [^1] In [aig], the term “strongly admissible sequence” is used for the bi-infinite sequence $^{\infty}s(t)^{\infty}$ in the sense of this paper, whereas here we use this name for its single block $s(t)$. . For $t=\frac{0}{1}$ or $t=\frac{1}{0}$, set $s(\frac{0}{1})=s(\frac{1}{0})=(1,1)$. For $t\in(0,\infty)$, we define $s(t)$ as follows:
  • (1)
    Set the orientation of $\overline{L_{t}}$ from left to right, and arrange the signs assigned to triangles by the triangle-crossing rule in the order in which $\overline{L_{t}}$ intersects them.
  • (2)
    Construct a sequence of integers $(a_{1},\dots,a_{\ell})$ from the counts of consecutive occurrences of the same sign in the sequence from (1). This sequence is called the strongly admissible sequence associated with $t$, denoted $s(t)$.
Example 2.14.
Let $t=\frac{2}{5}$. From Figure 7, the sign sequence is
and therefore
$s(\tfrac{2}{5})=(2,1,1,1,1,2,2,1,1,2).$
Figure 7. Signs of triangles intersecting $\overline{L_{t}}$ for $t=\frac{2}{5}$
Similarly, for $t=\frac{3}{2}$, Figure 8 shows the sign sequence
so that
$s(\tfrac{3}{2})=(2,2,2,2,1,1).$
Figure 8. Signs of triangles intersecting $\overline{L_{t}}$ for $t=\frac{3}{2}$
Remark 2.15.
We note that the following:
  • (1)
    If $t\in[1,\infty]$, then $c_{t}$ consists of letters $q$ and $r$, and $s(t)$ is obtained from $c_{t}$ by replacing each $q$ with “2,22,2” and each $r$ with “1,11,1”. In [aig, bombieri], $s(t)$ is constructed from $c_{t}$ by using this replacement.
  • (2)
    If $t\in(1,\infty)$, $s(t)$ has the form $(2,2,a_{3},\dots,a_{n-3},1,1)$. Moreover, in this case,
    $s\left(\tfrac{1}{t}\right)=(2,1,1,a_{3},\dots,a_{n-3},2).$
For $a\in\mathbb{Z}_{\geq 1}$, the matrix $\begin{bmatrix}a&1\\ 1&0\end{bmatrix}$ is called a continued fraction matrix. For any finite integer sequence $S$, define $CF_{S}$ as the matrix product obtained by substituting each $a\in S$ with the corresponding continued fraction matrix. The following theorem means that each Cohn matrix can be decomposed into continued fraction matrices.
Theorem 2.16.
For any irreducible fraction $t\in(0,\infty]$, we have
$C_{t}=CF_{s(t)}.$
Proof of Theorem 2.16 for $t\in[1,\infty]$.
Since
$Q=\begin{bmatrix}5&2\\ 2&1\end{bmatrix}=\begin{bmatrix}2&1\\ 1&0\end{bmatrix}^{2},\quad R=\begin{bmatrix}2&1\\ 1&1\end{bmatrix}=\begin{bmatrix}1&1\\ 1&0\end{bmatrix}^{2},$
and using Remark 2.15 (1), the theorem follows. ∎
We now consider the case $t\in(0,1)$.
Lemma 2.17.
For any irreducible fraction $t\in(1,\infty)$,
$CF_{s(\frac{1}{t})}=\begin{bmatrix}2n_{t}+u_{t}&\dfrac{2n_{t}^{2}-n_{t}u_{t}-u_{t}^{2}-1}{n_{t}}\\ n_{t}&n_{t}-u_{t}\end{bmatrix}$.
Proof.
By Remark 2.15 (2), $s(\frac{1}{t})$ is obtained from $s(t)$ by replacing the initial $2,2$ with $2,1,1$ and the final $1,1$ with $2$. The lemma follows from direct matrix computation (see also [bombieri2]*Corollary 24). ∎
Proof of Theorem 2.16 for $t\in(0,1)$.
By Theorem 2.11 and Lemmas 2.17 and 2.6, the $(2,1)$- and $(2,2)$-entries of $C_{t}$ and $CF_{s(t)}$ coincide. Since $\det(C_{t})=\det(CF_{s(t)})=1$ and $\mathrm{tr}(C_{t})=\mathrm{tr}(CF_{s(t)})=3n_{t}$, the claim follows. ∎
Remark 2.18.
In the proof of Theorem 2.16, we rely on the correspondence between the Cohn (Christoffel) word $c_{t}$ and the strongly admissible sequence $s(t)$. Later, we will prove this theorem in another way within a more general framework.
2.4. Markov-Lagrange Spectrum and Markov Spectrum
In this section, we will introduce the discrete Markov spectrum, which is a part of the Markov (and Markov-Lagrange) spectrum. We begin with a general theory of Markov-Lagrange spectrum and Markov spectrum.
Definition 2.19.
Let $\alpha\in\mathbb{R}$. Then $\mathcal{L}(\alpha)$ is the least upper bound of numbers $L$ such that
$\left|\alpha-\frac{p}{q}\right|<\frac{1}{Lq^{2}}$
holds for infinitely many rational approximations $\frac{p}{q}$, if this upper bound exists; otherwise, we define $\mathcal{L}(\alpha)=\infty$.
We define
$\mathcal{L}:=\{\mathcal{L}(\alpha)\mid\alpha\in\mathbb{R}\setminus\mathbb{Q}\},$
and call it the Markov-Lagrange spectrum.
If $\alpha$ is represented by a purely periodic regular continued fraction, the following calculation method of the Markov-Lagrange spectrum is known.
For any (possibly two-sided infinite) integer sequence $S$, let $S^{\ast}$ denote the reversed sequence. Let $P$ and $Q$ be right-sided infinite integer sequences. For the sequence $P=(a_{1},a_{2},\dots)$, define the regular continued fraction
$P]:=[a_{1},a_{2},\dots]=a_{1}+\frac{1}{a_{2}+\frac{1}{\ddots}}.$
For convergent cases, set
$\ell(P^{\ast}|Q):=\frac{1}{[P]}+[Q].$
For a finite positive integer sequence $T$, define the periodic sequences
$^{\infty}T^{\infty}:=\cdots TTTT\cdots,\quad^{\infty}T:=\cdots TT,\quad T^{\infty}:=TT\cdots.$
Set
$L(T):=\max_{{}^{\infty}T^{\infty}=P^{\ast}Q}\ell(P^{\ast}|Q).$
We have the following theorem [^2] In [hurwitz], a more general formula is given for computing $\mathcal{L}(\alpha)$ for an arbitrary irrational number $\alpha$ using its continued fraction expansion and $\lim\sup$. By applying this formula to the case of quadratic irrationals, one obtains Theorem 2.20 stated in this section. For a detailed discussion of this reformulation, see [bombieri]. :
Theorem 2.20 ([hurwitz]).
For any finite positive integer sequence $T$, we set $\alpha:=[T^{\infty}]$. We have
$\mathcal{L}(\alpha)=L(T).$
By Theorem 2.20, the value of $\mathcal{L}([T^{\infty}])$ can be computed by choosing a partition of $^{\infty}T^{\infty}$ that maximizes $\ell(P^{\ast}|Q)$.
Next, we will introduce the Markov spectrum.
Definition 2.21.
Let $Q(x,y)=ax^{2}+bxy+cy^{2}$ be a real indeterminate quadratic form. Then we define
$\mathcal{M}(Q):=\sup_{(x,y)\in\mathbb{Z}^{2}\setminus\{0\}}\frac{\sqrt{D}}{|Q(x,y)|},$
where $D=b^{2}-4ac$, if the supremum exists; otherwise, we define $\mathcal{M}(Q)=\infty$.
We define
$\mathcal{M}:=\{\mathcal{M}(Q)\mid Q\colon\text{real indefinite quadratic form}\},$
and call it the Markov spectrum.
For a finite positive integer sequence $T$, we set $\alpha:=[T^{\infty}]$ and $\alpha^{\prime}:=-[0,(T^{\ast})^{\infty}]$. Then, $\alpha^{\prime}$ is the quadratic conjugate of $\alpha$. We set
$Q_{T}=(x-\alpha y)(x-\alpha^{\prime}y).$
The quadratic form $Q_{T}$ has rational coefficients. We have the following [^3] In [perron2], a more general formula is given for computing $\mathcal{M}(Q)$ for arbitrary irreducible quadratic form $Q$ using its continued fraction expansion and $\sup$. By applying this formula to the case of quadratic irrationals, one obtains Theorem 2.22 stated in this section. For a detailed discussion of this reformulation, see [classical-dynamical]*Appendix C.:
Theorem 2.22 ([perron2]).
For any finite integer sequence $T$, we have
$\mathcal{M}(Q_{T})=L(T).$
In particular, we have $\mathcal{L}([T^{\infty}])=\mathcal{M}(Q_{T})$.
When $T$ is a strongly admissible sequence $s(t)$, the next theorem indicates that this maximizing partition can be explicitly determined.
Theorem 2.23 ([mar1, mar2]).
For any irreducible fraction $t\in[0,\infty]$,
  • (1)
    $s(t)^{\infty}]\in\mathbb{Q}[\sqrt{9n_{t}^{2}-4}]$, and
  • (2)
    we have
    $L(s(t))=\ell(^{\infty}s(t)|s(t)^{\infty})=\ell\left({}^{\infty}s\left(\tfrac{1}{t}\right)\,\middle|\,s\left(\tfrac{1}{t}\right)^{\infty}\right),$
    and this value is
    $\frac{\sqrt{9n_{t}^{2}-4}}{n_{t}}.$
Here, the value $9n_{t}^{2}=(3n_{t})^{2}$ is determined by the trace of $C_{t}$. Theorem 2.23 is proved within a more general framework in Section 6. The following corollary follows from Theorems 2.20, 2.22 and 2.23 immediately:
Corollary 2.24.
For any strongly admissible sequence $s(t)$, we have
$\mathcal{L}([s(t)^{\infty}])=\mathcal{M}(Q_{s(t)})=\frac{\sqrt{9n_{t}^{2}-4}}{n_{t}}.$
We set
$\mathcal{M}_{d}:=\left\{\frac{\sqrt{9n^{2}-4}}{n}\ \middle|\ n:\text{Markov number}\right\},$
called the discrete Markov spectrum.
There is an important characterization of $\mathcal{M}_{d}$. Let
$\mathcal{L}_{<3}$ $:=\{\mathcal{L}(\alpha)\mid\alpha\in\mathbb{R}\setminus\mathbb{Q}\text{ and }\mathcal{L}(\alpha)<3\},$
$\mathcal{M}_{<3}$ $:=\{\mathcal{M}(Q)\mid Q:\text{real indefinite quadratic form and }\mathcal{M}(Q)<3\}.$
Then, by Corollary 2.24 and the fact that $\frac{\sqrt{9n^{2}-4}}{n}<3$, we have $\mathcal{M}_{d}\subset\mathcal{L}_{<3}$ and $\mathcal{M}_{d}\subset\mathcal{M}_{<3}$. Moreover, the converse inclusion also holds, that is:
Theorem 2.25 ([mar1, mar2, hurwitz]).
We have $\mathcal{M}_{d}=\mathcal{L}_{<3}=\mathcal{M}_{<3}$.
3. Generalized Markov Equation and Generalized Markov Numbers
From this section, we consider the generalization of Markov equation. More precisely, we consider the $(k_{1},k_{2},k_{3})$-generalized Markov equation
$x^{2}+y^{2}+z^{2}+k_{1}yz+k_{2}zx+k_{3}xy=(3+k_{1}+k_{2}+k_{3})xyz,$
where $k_{1},k_{2},k_{3}\in\mathbb{Z}_{\geq 0}$. We abbreviate this equation as the $(k_{1},k_{2},k_{3})$-GM equation.
A permutation of a positive integer solution to this equation is called a $(k_{1},k_{2},k_{3})$-GM triple, and a positive integer appearing in some $(k_{1},k_{2},k_{3})$-GM triple is called a $(k_{1},k_{2},k_{3})$-GM number. Note that the definition of Markov triples includes all permutations of a solution.
We introduce the $(k_{1},k_{2},k_{3},\sigma)$-generalized Markov tree (abbreviated $(k_{1},k_{2},k_{3},\sigma)$-GM tree). Let $\mathfrak{S}_{3}$ be the symmetric group on 3 elements, acting on $\{1,2,3\}$ from the left.
Define the $(k_{1},k_{2},k_{3},\sigma)$-GM tree $\mathrm{M}\mathbb{T}(k_{1},k_{2},k_{3},\sigma)$ for $\sigma\in\mathfrak{S}_{3}$ as follows:
  • (1)
    The root vertex is
    $\big((1,\sigma(1)),(k_{\sigma(2)}+2,\sigma(2)),(1,\sigma(3))\big).$
  • (2)
    Every vertex $((a,h),(b,i),(c,j))$ has the following two children:
    ${((a,h),(b,i),(c,j))}$${\left((a,h),\left(\frac{a^{2}+k_{j}ab+b^{2}}{c},j\right),(b,i)\right)}$${\left((b,i),\left(\frac{b^{2}+k_{h}bc+c^{2}}{a},h\right),(c,j)\right).}$
Example 3.1.
The first few vertices of $\mathrm{M}\mathbb{T}(1,2,0,\mathrm{id})$ are as follows:
${((1,1),(4,2),(1,3))}$${((4,2),(21,1),(1,3))}$${((1,1),(17,3),(4,2))}$${((21,1),(121,2),(1,3))}$${((4,2),(457,3),(21,1))}$${((17,3),(373,1),(4,2))}$${((1,1),(81,2),(17,3))}$${((121,2),(703,1),(1,3))\cdots}$${((21,1),(15082,3),(121,2))\cdots}$${((457,0),(57121,2),(21,1))\cdots}$${((4,2),(10033,1),(457,3))\cdots}$${((373,1),(8185,3),(4,2))\cdots}$${((17,3),(38025,2),(373,1))\cdots}$${((81,2),(8227,1),(17,3))\cdots}$${((1,1),(386,3),(81,2))\cdots}$
For any $\tau\in\mathfrak{S}_{3},$ we set $\tau(x_{1},x_{2},x_{3}):=(x_{\tau(1)},x_{\tau(2)},x_{\tau(3)})$. For any $\sigma\in\mathfrak{S}_{3}$, we denote by $\sigma^{\ast}\in\mathfrak{S}_{3}$ an element satisfying $\sigma(2)=\sigma^{\ast}(2)$. We have the following theorem.
Theorem 3.2.
Let $(k_{1},k_{2},k_{3})$-GM triple $(a,b,c)$ with $b>\max\{a,c\}$, and assume $\tau(a,b,c)$ is a solution to the $(k_{1},k_{2},k_{3})$-GM equation for $\tau\in\mathfrak{S}_{3}$. Then there exist a unique pair $\{\sigma,\sigma^{*}\}\subset\mathfrak{S}_{3}$ and a unique pair $(v,v^{\ast})$ such that $v\in\mathrm{M}\mathbb{T}(k_{1},k_{2},k_{3},\sigma)$ and $v^{\ast}\in\mathrm{M}\mathbb{T}(k_{1},k_{2},k_{3},\sigma^{\ast})$, and $v$ and $v^{\ast}$ are permutations of $((a,\tau^{-1}(1)),(b,\tau^{-1}(2)),(c,\tau^{-1}(3)))$. Furthermore, the position of $v^{\ast}$ in $\mathrm{M}\mathbb{T}(k_{1},k_{2},k_{3},\sigma^{\ast})$ is symmetric with respect to the center to the position of $v$ in $\mathrm{M}\mathbb{T}(k_{1},k_{2},k_{3},\sigma)$.
Proof.
It follows from [bana-gyo]*Proposition 3.8 and [gyomatsu]*Theorem 1. ∎
Each pair within the triples forming the vertices of $\mathrm{M}\mathbb{T}(k_{1},k_{2},k_{3},\sigma)$ has its first component representing a $(k_{1},k_{2},k_{3})$-GM number. Moreover, for each triple, the three numbers given by the first components form a solution to the $(k_{1},k_{2},k_{3})$-GM equation when placed in the positions indicated by the second components of the corresponding pairs. These pairs are called $(k_{1},k_{2},k_{3},\sigma)$-GM-number-position pairs.
Remark 3.3.
Note that the following two facts:
  • (1)
    If $k_{1}=k_{2}=k_{3}$, then $(k_{1},k_{2},k_{3},\sigma)$-GM tree is the same tree up to position element. We often omit $\sigma$ in this case.
  • (2)
    If $b=\max\{a,c\}$, then $(a,b,c)=(1,1,1)$ ([gyomatsu]*Lemma 4), therefore the set
    $\{v\in\mathrm{M}\mathbb{T}(k_{1},k_{2},k_{3},\sigma)\mid\sigma\in\mathfrak{S}_{3}\}$
    contains essentially all $(k_{1},k_{2},k_{3})$-GM triples except $(1,1,1)$.
The following proposition is an analogue of Proposition 2.2.
Proposition 3.4 ([gyomatsu]*Corollary 8).
For any $(k_{1},k_{2},k_{3})$-GM triple $(a,b,c)$, all pairs in $a,b,c$ are relatively prime.
Now, fix $(k_{1},k_{2},k_{3})\in\mathbb{Z}_{\geq 0}^{3}$ and $\sigma\in\mathfrak{S}_{3}$.
Abusing notation as in the previous section, we denote by $(n_{t},i_{t})$ the $(k_{1},k_{2},k_{3},\sigma)$-GM-number-position pair in $\mathrm{M}\mathbb{T}(k_{1},k_{2},k_{3},\sigma)$ located at the position of the irreducible fraction $t\in[0,\infty]$ in $\mathrm{F}\mathbb{T}$.
For any $(k_{1},k_{2},k_{3})$-GM triple $(n_{r},n_{t},n_{s})$ in $\mathrm{M}\mathbb{T}(k_{1},k_{2},k_{3},\sigma)$, consider an integer $u_{t}$ satisfying:
$\begin{cases}n_{r}u_{t}\equiv n_{s}\pmod{n_{t}},\\ 0<u_{t}<n_{t}.\end{cases}$
The integer $u_{t}$ is uniquely determined since $n_{r},n_{t},n_{s}$ are pairwise relatively prime. Moreover, $u_{t}$ depends only on $t$, since $(r,t,s)$ is determined by $t$ from Proposition 2.5 (2).
For $t\in(0,\infty)$, we define $u_{t}$ as above, and set
$u_{\frac{0}{1}}=-k_{\sigma(1)},\quad u_{\frac{1}{0}}=1.$
The number $u_{t}$ is called the characteristic number of $t$.
Remark 3.5.
Let $(n^{\ast}_{\frac{1}{t}},i^{*}_{\frac{1}{t}})$ be a $(k_{1},k_{2},k_{3},\sigma^{\ast})$-GM number-position pair associated with $t$. Then, by Theorem 3.2, we have
$(n_{t},i_{t})=(n_{\frac{1}{t}}^{\ast},i^{\ast}_{\frac{1}{t}}).$
We abbreviate $k_{i_{t}}$ as $k_{t}$. The following lemma is generalization of Lemma 2.6:
Lemma 3.6.
For any irreducible fraction $t\in(0,\infty)$,
$u_{\frac{1}{t}}=n_{t}-u^{\ast}_{t}-k_{t}.$
The proof of Lemma 3.6 can be seen [gyoda-maruyama-sato]*Proposition 7.22 (2) for the case where $k_{1}=k_{2}=k_{3}$. We can prove the general case in the same way. As in the case of Lemma 2.6, replace $v_{t}^{+}$ in [gyoda-maruyama-sato] with $u^{*}_{\frac{1}{t}}$ and take the duality $(\ )^{\ast}$.
4. Generalized Cohn Matrices and their Decompositions
In this section, we will introduce analogue of Cohn matrices corresponding to GM numbers, and give continued fraction matrix decompositions of them.
4.1. Generalized Cohn Matrices
Fix $k_{1},k_{2},k_{3}\in\mathbb{Z}_{\geq 0}$ and $\sigma\in\mathfrak{S}_{3}$. We set
$C_{\frac{0}{1}}$ $=\begin{bmatrix}3+k_{1}+k_{2}+k_{3}&-(3+k_{1}+k_{2}+k_{3})k_{\sigma(1)}-1\\ 1&-k_{\sigma(1)}\end{bmatrix}$,
$C_{\frac{1}{1}}$ $=\begin{bmatrix}(3+k_{1}+k_{2}+k_{3})(k_{\sigma(2)}+2)-k_{\sigma(2)}-1&2+k_{1}+k_{2}+k_{3}\\ k_{\sigma(2)}+2&1\end{bmatrix}$,
$C_{\frac{1}{0}}$ $=\begin{bmatrix}2+k_{1}+k_{2}+k_{3}-k_{\sigma(3)}&1+k_{1}+k_{2}+k_{3}-k_{\sigma(3)}\\ 1&1\end{bmatrix}$.
For any $(r,t,s)\in\mathrm{F}\mathbb{T}$, we set inductively
$C_{r\oplus t}:=C_{r}C_{t}-D_{s},\quad C_{t\oplus s}=C_{t}C_{s}-D_{r},$
where
$D_{r}=\begin{bmatrix}k_{r}&k_{r}(3+k_{1}+k_{2}+k_{3})\\ 0&k_{r}\end{bmatrix}$.
These matrices coincide with ordinary Cohn matrices when $k_{1}=k_{2}=k_{3}=0$. The matrix $C_{t}$ is called a $(k_{1},k_{2},k_{3},\sigma)$-generalized Cohn matrix. Define the $(k_{1},k_{2},k_{3},\sigma)$-generalized Cohn tree:
$\mathrm{Co}\mathbb{T}(k_{1},k_{2},k_{3},\sigma):=\mathrm{F}\mathbb{T}\big|_{(r,t,s)\mapsto(C_{r},C_{t},C_{s})}.$
Example 4.1.
The first few vertices in $\mathrm{Co}\mathbb{T}(1,2,0,\mathrm{id})$ are:
${\left(\begin{bmatrix}6&-7\\ 1&-1\end{bmatrix},\begin{bmatrix}21&5\\ 4&1\end{bmatrix},\begin{bmatrix}5&4\\ 1&1\end{bmatrix}\right)}$${\left(\begin{bmatrix}6&-7\\ 1&-1\end{bmatrix},\begin{bmatrix}98&23\\ 17&4\end{bmatrix},\begin{bmatrix}21&5\\ 4&1\end{bmatrix}\right)}$${\left(\begin{bmatrix}21&5\\ 4&1\end{bmatrix},\begin{bmatrix}109&83\\ 21&16\end{bmatrix},\begin{bmatrix}5&4\\ 1&1\end{bmatrix}\right)}$${\left(\begin{bmatrix}98&23\\ 17&4\end{bmatrix},\begin{bmatrix}2149&507\\ 373&88\end{bmatrix},\begin{bmatrix}21&5\\ 4&1\end{bmatrix}\right)}$${\left(\begin{bmatrix}6&-7\\ 1&-1\end{bmatrix},\begin{bmatrix}367&98\\ 81&17\end{bmatrix},\begin{bmatrix}98&23\\ 17&4\end{bmatrix}\right)}$${\left(\begin{bmatrix}109&83\\ 16&21\end{bmatrix},\begin{bmatrix}626&507\\ 121&98\end{bmatrix},\begin{bmatrix}5&4\\ 1&1\end{bmatrix}\right)}$${\left(\begin{bmatrix}21&5\\ 4&1\end{bmatrix},\begin{bmatrix}2394&1823\\ 457&348\end{bmatrix},\begin{bmatrix}109&83\\ 21&16\end{bmatrix}\right)}$
Theorem 4.2.
For any irreducible fraction $t\in[0,\infty]$, we have
$C_{t}=\begin{bmatrix}(3+k_{1}+k_{2}+k_{3})n_{t}-k_{t}-u_{t}&\dfrac{(3+k_{1}+k_{2}+k_{3})n_{t}u_{t}-k_{t}u_{t}-u_{t}^{2}-1}{n_{t}}\\ n_{t}&u_{t}\end{bmatrix}$.
In particular, we have $\det(C_{t})=1$ and $\mathrm{tr}(C_{t})=(3+k_{1}+k_{2}+k_{3})n_{t}-k_{t}$.
In [gyo-maru]*Lemma 4.5, Theorem 4.2 is treated for $k_{1}=k_{2}=k_{3}$ and $t\in[0,1]$. The general case is proved similarly.
4.2. Continued Fraction Matrix Decomposition of Generalized Cohn Matrices
As in Section 2, we decompose $C_{t}$ into a product of fundamental matrices using geometric information. For any irreducible fraction $t\in[0,\infty]$, we construct a generalized strongly admissible sequence $s(t)$ by the triangle-crossing rule introduced in Section 2 and the following edge-crossing rule.
Definition 4.3.
We fix $(k_{1},k_{2},k_{3})\in\mathbb{Z}_{\geq 0}^{3}$ and $\sigma\in\mathfrak{S}_{3}$. Let $\gamma$ be an oriented curve segment on $\widetilde{\mathbb{R}^{2}}$. For each edge in a triangle intersecting the interior of $\gamma$, assign signs as follows:
  • (i)
    Assign $k_{\sigma(1)}$ (resp., $k_{\sigma(2)}$, $k_{\sigma(3)}$) minus signs (-) to each horizontal (resp., diagonal, vertical) edge whose midpoint is not strictly on the right side of $\gamma$ (see Figure 9).
    $\mathbf{<}$
    $\mathbf{>}$
    $\mathbf{<}$
    $\mathbf{>}$
    $\mathbf{<}$
    $\mathbf{>}$
    Figure 9. Edges with -
  • (ii)
    Assign $k_{\sigma(1)}$ (resp., $k_{\sigma(2)}$, $k_{\sigma(3)}$) plus signs (++) to each horizontal (resp., diagonal, vertical) edge whose midpoint is on the right side of $\gamma$ (see Figure 10).
    $\mathbf{>}$
    $\mathbf{<}$
    $\mathbf{>}$
    $\mathbf{<}$
    $\mathbf{>}$
    $\mathbf{<}$
    Figure 10. Edges with ++
This rule is called the edge-crossing rule of $\gamma$ for $(k_{1},k_{2},k_{3},\sigma)$.
We define a $(k_{1},k_{2},k_{3},\sigma)$-generalized strongly admissible sequence $s(t)$. First, we set
$s(\tfrac{0}{1})=(1+k_{\sigma(2)}+k_{\sigma(3)},1),\quad s(\tfrac{1}{0})=(1+k_{\sigma(1)}+k_{\sigma(2)},1)$
For any irreducible fraction $t\in(0,\infty)$, we define $s(t)$ as follows:
  • (1)
    Orient $\overline{L_{t}}$ from left to right. For any irreducible fraction $t\in(0,\infty)$, arrange the signs assigned to the triangles by the triangle-crossing and edge-crossing rules in the order in which $\overline{L}_{t}$ passes through them.
  • (2)
    Construct the integer sequence $(a_{1},\dots,a_{\ell})$ from the numbers of consecutive occurrences of the same sign in (1). It is called the generalized strongly admissible sequence associated with $t$, denoted $s(t)$.
Example 4.4.
Let $(k_{1},k_{2},k_{3},\sigma)=(1,2,0,\textrm{id})$ and $t=\frac{2}{5}$. Figure 11 shows the signs along $\overline{L}_{t}$, giving
$s(\tfrac{2}{5})=(5,1,3,3,1,5,4,1,3,4).$
Figure 11. Signs of triangles and edges intersecting with $\overline{L}_{t}$ for $t=\frac{2}{5}$
The right endpoint of $\overline{L_{t}}$ lies within the upper-rightmost edge in $\widetilde{\mathbb{R}^{2}}$, therefore no sign is assigned to this edge. Although the unique horizontal edge intersecting $\overline{L_{t}}$ appears to meet it at its midpoint, in fact, since $\overline{L_{t}}$ is defined by shifting $L_{t}$ slightly to the left, the intersection point lies slightly to the left of the midpoint. Hence, note that the corresponding sign is ++.
Remark 4.5.
For $s(t)=(a_{1},\dots,a_{n})$:
  • (1)
    $a_{1}=2+k_{1}+k_{2}+k_{3}$.
  • (2)
    If $t=\frac{1}{1}$, then $s(t)=(2+k_{1}+k_{2}+k_{3},2+k_{\sigma(2)}).$
  • (3)
    If $t\in(0,1)$, then $a_{2}=1$, and $a_{n}\neq 1$. If $t\in(1,\infty)$, then $a_{2}\neq 1$, and $a_{n}=1$.
  • (4)
    If $t=\frac{1}{2}$, then $n=4$ and $a_{4}=a_{3}+1+k_{t}$. If $t=\frac{2}{1}$, then $n=4$ and $a_{2}=a_{3}+1+k_{t}$.
  • (5)
    If $t\in(0,\frac{1}{2})\cup(\frac{1}{2},1)$, then $a_{3}+1=a_{n}$, $a_{3+i}=a_{n-i}$ for $i=1,2,\dots,\frac{n}{2}-3$. Moreover, we have $a_{\frac{n}{2}+1}=a_{\frac{n}{2}+2}+(-1)^{\frac{n}{2}+1}k_{t}$.
  • (6)
    If $t\in(1,\frac{2}{1})\cup(\frac{2}{1},\infty)$, then $a_{2}=a_{n-1}+1$, $a_{2+i}=a_{n-i-1}$ for $i=1,2,\dots,\frac{n}{2}-3$. Moreover, we have $a_{\frac{n}{2}}=a_{\frac{n}{2}+1}+(-1)^{\frac{n}{2}}k_{t}$.
  • (7)
    For $t\in(0,\infty)$, denote by $s^{\ast}(\frac{1}{t})$ the $(k_{1},k_{2},k_{3},\sigma^{\ast})$-generalized strongly admissible sequence, expressed as $(a_{1},a_{n},a_{n-1},\dots,a_{2})$. Clearly, $^{\infty}(s^{*}(\tfrac{1}{t}))^{\infty}$ is the reverse sequence of $^{\infty}s(t)^{\infty}$. This fact also holds for $t=\frac{0}{1}$ and $\frac{1}{0}$.
We will prove the following theorem:
Theorem 4.6.
For any irreducible fraction $t\in(0,\infty]$, we have $C_{t}=CF_{s(t)}$.
To prove Theorem 4.6, we introduce the snake graph. Assume $a_{i}\in\mathbb{Z}_{\geq 1}$. First, recall the snake graph associated with a continued fraction $a_{1},\dots,a_{\ell}]$ according to [cs18]:
For a continued fraction $a_{1},\dots,a_{\ell}]$ with $(\ell,a_{1})\neq(1,1)$:
  1. (1)
    Arrange $(a_{1}+\cdots+a_{\ell})$ signs: the first $a_{1}$ are -, the next $a_{2}$ are ++, alternating between - and ++.
  2. (2)
    Remove the first and last signs to obtain a tuple $S$ of length $(a_{1}+\cdots+a_{\ell}-2)$.
  3. (3)
    Using $S$, place tiles as in Figure 12, satisfying:
    • The first (leftmost) tile is the left one in Figure 12.
    • Each new tile is placed to the right or above the preceding tile.
    • Signs on adjoining edges coincide.
    • $S$ coincides with the sequence of signs on adjoining parts from left to right.
This graph is the snake graph, denoted $\mathcal{G}[a_{1},\dots,a_{\ell}]$. For the empty continued fraction [][\ ], set $\mathcal{G}[\ ]=\emptyset$. For $[1]$, set $\mathcal{G}[1]$ as a line segment. $\mathcal{G}[2]$ is a single tile.
Figure 12. Signed tiles
Example 4.7.
For $[2,4,2,1]$, the snake graph is shown in Figure 13. Indeed, the signs located inside the connected tiles are arranged from the leftmost to the rightmost as follows: there are (212-1) consecutive “-” signs, followed by 4 “++” signs, then 2 “-” signs, and finally (11)(1-1) (therefore, no) consecutive “++” signs.
Figure 13. Snake graph associated with $[2,4,2,1]$
Remark 4.8.
The sign on the right edge in each tile is different from one on the upper edge. Therefore, for a continued fraction $a_{1},\dots,a_{\ell}]$, there is a unique snake graph associated with $a_{1},\dots,a_{\ell}]$.
Let $G$ be an undirected graph. We recall that a subset $P$ of the edge set of $G$ is called a perfect matching of $G$ if each vertex of $G$ is incident to exactly one edge in $P$. We denote by $m(\mathcal{G}[a_{1},\dots,a_{\ell}])$ the number of perfect matchings of $\mathcal{G}[a_{1},\dots,a_{\ell}]$. We set $m(\mathcal{G}[\ ])=1$.
Example 4.9.
By an induction argument, we have $m(\mathcal{G}[n])=n$. The Figure 14 is the list of perfect matchings of $\mathcal{G}[5]$.
Figure 14. List of perfect matchings of $\mathcal{G}[5]$
The following proposition follows from the fact that the number of perfect matchings of a snake graph is invariant under congruent transformations of the graph.
Proposition 4.10.
For any integer sequence $(a_{1},\dots,a_{n})$, we have
$m(\mathcal{G}[a_{1},\dots,a_{n}])=m(\mathcal{G}[a_{n},\dots,a_{1}]).$
In [cs18], the authors establish the following relation between a continued fraction and its associated snake graph:
Theorem 4.11 ([cs18]*Theorem 3.4).
The following equality holds:
$a_{1},\dots,a_{n}]=\frac{m(\mathcal{G}[a_{1},\dots,a_{n}])}{m(\mathcal{G}[a_{2},\dots,a_{n}])}.$
We note that Theorem 4.11 also covers the case in which the denominator on the right-hand side is $m(\mathcal{G}[\ ])$.
If $a_{1},\dots,a_{n}]=\frac{p_{n}}{q_{n}}$, then the corresponding product of continued fraction matrices is given by (see [bombieri]*Section 2.1):
$\begin{bmatrix}a_{1}&1\\ 1&0\end{bmatrix}\begin{bmatrix}a_{2}&1\\ 1&0\end{bmatrix}\cdots\begin{bmatrix}a_{n}&1\\ 1&0\end{bmatrix}=\begin{bmatrix}p_{n}&p_{n-1}\\ q_{n}&q_{n-1}\end{bmatrix}$.
By Theorem 4.11, this can be restated as
(4.1) $\begin{bmatrix}a_{1}&1\\ 1&0\end{bmatrix}\begin{bmatrix}a_{2}&1\\ 1&0\end{bmatrix}\cdots\begin{bmatrix}a_{n}&1\\ 1&0\end{bmatrix}=\begin{bmatrix}m(\mathcal{G}[a_{1},\dots,a_{n}])&m(\mathcal{G}[a_{1},\dots,a_{n-1}])\\ m(\mathcal{G}[a_{2},\dots,a_{n}])&m(\mathcal{G}[a_{2},\dots,a_{n-1}])\end{bmatrix}$.
Hence, to prove Theorem 4.6, it suffices to show that for $s(t)=(a_{1},\dots,a_{n})$,
(4.2) $C_{t}=\begin{bmatrix}m(\mathcal{G}[a_{1},\dots,a_{n}])&m(\mathcal{G}[a_{1},\dots,a_{n-1}])\\ m(\mathcal{G}[a_{2},\dots,a_{n}])&m(\mathcal{G}[a_{2},\dots,a_{n-1}])\end{bmatrix}$
for any $t\in(0,\infty]$.
On the other hand, [gyoda-maruyama-sato]*Theorem 7.10 provides an expression for $C_{t}$ for $t\in(0,1)$. Adapting it to the notation of the present paper, we obtain the following:
Theorem 4.12.
Let $t\in(0,1)$ be an irreducible fraction. For $s(t)=(a_{1},\dots,a_{n})$, we have
[…]
Remark 4.13.
Note that when comparing Theorem 4.12 with [gyoda-maruyama-sato]*Theorem 7.10 (2), the positions of the matrix entries differ due to the difference in the convention of Cohn matrices described in Remark 2.10.
By using Theorem 4.12, we can describe $C_{t}$ for $t\in(1,\infty)$:
Theorem 4.14.
Let $t\in(1,\infty)$ be an irreducible fraction. For $s(t)=(a_{1},\dots,a_{n})$, we have
[…]
Remark 4.15.
The $(1,2)$ entry in (4.4) is omitted since it does not need to be proved, but it is actually
$(2+k_{1}+k_{2}+k_{3})\,m(\mathcal{G}[a_{n-1},\dots,a_{2}])+m(\mathcal{G}[a_{n-1},\dots,a_{3}]).$
To prove Theorem 4.14, we need the snake graph calculus. The following two lemmas are essentially proved by [banaian-sen]. In this paper, the statement is formulated in terms of the numerators of continued fractions, but by applying the Theorem 4.11 it can be rephrased as a statement about the number of perfect matchings.
Lemma 4.16 ([banaian-sen]*Lemma 3).
Let $n\geq 2$. For any integer sequence $(a_{1},\dots,a_{n})$, the following equality holds:
$m(\mathcal{G}[a_{1},\dots,a_{n}])=a_{n}m(\mathcal{G}[a_{1},\dots,a_{n-1}])+m(\mathcal{G}[a_{1},\dots,a_{n-2}]),$
where $m(\mathcal{G}[a_{1},a_{0}])=1$.
We call a sequence of the form $(a_{1},\dots,a_{\ell-1},\,a_{\ell}+k,\,a_{\ell},\,a_{\ell-1},\dots,a_{1})$ a semi-palindromic sequence, where $k\in\mathbb{Z}$.
Lemma 4.17.
For any semi-palindromic integer sequence $(a_{1},\dots,a_{\ell-1},a_{\ell}+k,a_{\ell},a_{\ell-1},\dots,a_{1})$, the following identity holds:
$m(\mathcal{G}[a_{2},\dots,a_{\ell-1},a_{\ell}+k,a_{\ell},a_{\ell-1},\dots,a_{1}])=m(\mathcal{G}[a_{1},\dots,a_{\ell-1},a_{\ell}+k,a_{\ell},a_{\ell-1},\dots,a_{2}])+(-1)^{\ell}k.$
Moreover, this identity is equivalent to
$m(\mathcal{G}[a_{2},\dots,a_{\ell-1},a_{\ell},a_{\ell}+k,a_{\ell-1},\dots,a_{1}])=m(\mathcal{G}[a_{1},\dots,a_{\ell-1},a_{\ell},a_{\ell}+k,a_{\ell-1},\dots,a_{2}])+(-1)^{\ell+1}k.$
Proof.
For the first statement of Lemma 4.17, the case where $k=1$ is proved in [banaian-sen]*Lemma 5. We can prove other cases in the same way as [banaian-sen]*Lemma 5. The second statement follows from Proposition 4.10. ∎
We will prove Theorem 4.14.
Proof of Theorem 4.14.
Let $s^{\ast}(\tfrac{1}{t})=(b_{1},\dots,b_{\ell})$. Since $b_{2}=1$ by Remark 4.5 (3), we have
$m(\mathcal{G}[b_{\ell},\dots,b_{3}+1])=m(\mathcal{G}[b_{\ell},\dots,b_{2}]).$
By Theorem 4.2 and Theorem 4.12, it follows that
$n^{\ast}_{\frac{1}{t}}=m(\mathcal{G}[b_{\ell},\dots,b_{2}]),\quad u^{\ast}_{\frac{1}{t}}=m(\mathcal{G}[b_{\ell-1},\dots,b_{2}]).$
Moreover, by Remark 4.5 (7), we have $\ell=n$, $b_{1}=a_{1}$, and $b_{i}=a_{n-i+2}$ for $2\leq i\leq n$. We recall Remark 3.5 and Lemma 3.6, that is,
$n_{t}=n^{\ast}_{\frac{1}{t}},\qquad u_{t}=n_{t}-u^{\ast}_{\frac{1}{t}}-k_{t}.$
Thus, we have
(4.5) $n_{t}$ $=m(\mathcal{G}[a_{2},\dots,a_{n}]),$
(4.6) $u_{t}$ $=m(\mathcal{G}[a_{2},\dots,a_{n}])-m(\mathcal{G}[a_{3},\dots,a_{n}])-k_{t}.$
Since $a_{n}=1$ by Remark 4.5 (3), we obtain
(4.7) $m(\mathcal{G}[a_{3},\dots,a_{n}])=m(\mathcal{G}[a_{3},\dots,a_{n-1}+1]).$
Furthermore, since $(a_{2},\dots,a_{n-1}+1)$ is a semi-palindrome by Remark 4.5 (4) and (6), we obtain
[…]
from Lemma 4.17. Therefore, substituting (4.8) into (4.6), we have
$u_{t}=m(\mathcal{G}[a_{2},\dots,a_{n}])-m(\mathcal{G}[a_{2},\dots,a_{n-2}]).$
Since $a_{n}=1$ again, Lemma 4.16 yields
$m(\mathcal{G}[a_{2},\dots,a_{n}])=m(\mathcal{G}[a_{2},\dots,a_{n-1}])+m(\mathcal{G}[a_{2},\dots,a_{n-2}]).$
Hence, we conclude that
$u_{t}=m(\mathcal{G}[a_{n-1},\dots,a_{2}]).$
By Theorem 4.2, it follows that
$C_{t}=\begin{bmatrix}\ast&\ast\\ m(\mathcal{G}[a_{n},\dots,a_{2}])&m(\mathcal{G}[a_{n-1},\dots,a_{2}])\end{bmatrix}$.
Since $\mathrm{tr}(C_{t})=(3+k_{1}+k_{2}+k_{3})n_{t}-k_{t}$, using (4.8), the $(1,1)$-entry is
$(2+k_{1}+k_{2}+k_{3})\,m(\mathcal{G}[a_{n},\dots,a_{2}])+m(\mathcal{G}[a_{n},\dots,a_{3}]).$
To prove Theorem 4.6, we will introduce another relation between snake graphs:
Lemma 4.18.
Let $n\geq 4$. For any sequence $(a_{1},1,a_{3},\dots,a_{n})$, we have
$m(\mathcal{G}[a_{1},1,a_{3},\dots,a_{n}])=(a_{1}+1)m(\mathcal{G}[a_{3}+1,\dots,a_{n}])-m(\mathcal{G}[a_{4},\dots,a_{n-1}]),$
where $m(\mathcal{G}[a_{4},a_{3}])=1$.
The special case of Lemma 4.18 is proved by [gyoda-maruyama-sato]*(7.2). Lemma 4.18 can be proved in the same way.
Proof of Theorem 4.6.
The cases $t=\frac{1}{1}$ and $t=\frac{1}{0}$ can be checked directly. We focus on $t\in(0,1)$. Comparing (4.2) and (4.3), it suffices to show
(4.9) $m(\mathcal{G}[a_{1},\dots,a_{n}])$ $=(3+k_{1}+k_{2}+k_{3})m(\mathcal{G}[a_{n},\dots,a_{3}+1])-m(\mathcal{G}[a_{n},\dots,a_{4}]),$
(4.10) $m(\mathcal{G}[a_{1},\dots,a_{n-1}])$ $=(3+k_{1}+k_{2}+k_{3})m(\mathcal{G}[a_{n-1},\dots,a_{3}+1])-m(\mathcal{G}[a_{n-1},\dots,a_{4}]),$
(4.11) $m(\mathcal{G}[a_{2},\dots,a_{n}])$ $=m(\mathcal{G}[a_{n},\dots,a_{3}+1]),$
(4.12) $m(\mathcal{G}[a_{2},\dots,a_{n-1}])$ $=m(\mathcal{G}[a_{n-1},\dots,a_{3}+1]).$
Since $a_{2}=1$, we have
$m(\mathcal{G}[a_{n},\dots,a_{3}+1])=m(\mathcal{G}[a_{n},\dots,a_{2}]),\quad m(\mathcal{G}[a_{n-1},\dots,a_{3}+1])=m(\mathcal{G}[a_{n-1},\dots,a_{2}]).$
Since $\mathcal{G}[a_{n},\dots,a_{2}]$ and $\mathcal{G}[a_{2},\dots,a_{n}]$ are mirror images of each other as graphs, the values $m(\mathcal{G}[a_{n},\dots,a_{2}])$ and $m(\mathcal{G}[a_{2},\dots,a_{n}])$ are equal. Therefore, (4.11) and (4.12) hold.
By Remark 4.5 (1), Equation (4.9) can be rewritten as
$m(\mathcal{G}[2+k_{1}+k_{2}+k_{3},1,a_{3},\dots,a_{n}])=(3+k_{1}+k_{2}+k_{3})m(\mathcal{G}[a_{n},\dots,a_{3}+1])-m(\mathcal{G}[a_{n-1},\dots,a_{4}]).$
This is the special case $a_{1}=2+k_{1}+k_{2}+k_{3}$ of Lemma 4.18. We can prove (4.10) in the same way. Next, we will prove the statement for $t\in(1,\infty)$. Since $\det(CF_{s(t)})=1$ and $\det C_{t}=1$ by Theorem 4.2, it suffices to show the coincidence of (1,1)(1,1)-,$(2,1)$-, and(2,2)(2,2)- entries. Comparing (4.2) and (4.4), it suffices to show
(4.13) $m(\mathcal{G}[a_{1},\dots,a_{n}])$ $=(2+k_{1}+k_{2}+k_{3})m(\mathcal{G}[a_{n},\dots,a_{2}])+m(\mathcal{G}[a_{n},\dots,a_{3}]),$
(4.14) $m(\mathcal{G}[a_{2},\dots,a_{n}])$ $=m(\mathcal{G}[a_{n},\dots,a_{2}]),$
(4.15) $m(\mathcal{G}[a_{2},\dots,a_{n-1}])$ $=m(\mathcal{G}[a_{n-1},\dots,a_{2}]).$
The equation (4.13) follows from Lemma 4.16, and (4.14) and (4.15) are trivial. ∎
Remark 4.19.
From definition, we can find that $C_{\frac{0}{1}}\neq CF_{s(\frac{0}{1})}$. However, the main theorem (Theorem 6.1) is correct even if $t=\frac{0}{1}$. We prove this theorem by direct computation only in the case $t=\frac{0}{1}$.
5. Generalized Markov Length and Generalized Markov Distance
In this section, we will introduce the key theorem (Theorem 5.7) to generalize the discrete Markov spectrum. We will begin with the definition of a generalized arc.
Definition 5.1.
Let $\gamma$ be a curve segment in $\widetilde{\mathbb{R}^{2}}$. If both endpoints of $\gamma$ are lattice points and $\gamma$ does not intersect the same edge of $\widetilde{\mathbb{R}^{2}}$ twice in succession, then $\gamma$ is called a generalized arc.
To define the generalized Markov length (GM length for short), we introduce a sign rule for generalized arcs.
Definition 5.2.
Let $\gamma$ be an oriented generalized arc on $\widetilde{\mathbb{R}^{2}}$. We assign a sign ()(-) or $(+)$ to the triangle one of whose angle coincides with the endpoint of $\gamma$, provided that $\gamma$ intersects the opposite edge to this angle (see Figure 15).
$\mathbf{<}$
$\mathbf{>}$
$\mathbf{>}$
$\mathbf{<}$
$\mathbf{>}$
$\mathbf{<}$
$\mathbf{>}$
$\mathbf{<}$
$\mathbf{<}$
$\mathbf{>}$
$\mathbf{<}$
$\mathbf{>}$
Figure 15. Right-angled triangles with - or ++ signs assigned according to the endpoint rule.
This procedure for assigning signs to triangles is called the endpoint rule of $\gamma$.
Remark 5.3.
The signs assigned by the endpoint rule do not affect the shape of the snake graph constructed from the entire sign sequence. From this viewpoint, we regard the signs ++ and - assigned by the endpoint rule as interchangeable.
We fix $(k_{1},k_{2},k_{3})\in\mathbb{Z}_{\geq 0}^{3}$ and $\sigma\in\mathfrak{S}_{3}$. Assume that $\gamma$ is a generalized arc. We define a sign sequence $s(\gamma)$ according to the following procedure:
  • (1)
    Arrange the signs assigned to triangles by the triangle-crossing, edge-crossing, and endpoint rules of $\gamma$ in the order that $\gamma$ passes through them.
  • (2)
    Construct a sequence of integers $(a_{1},\dots,a_{\ell})$ from the numbers of consecutive occurrences of the same sign in the sequence obtained in (1). This sequence is called the sign sequence associated with $\gamma$ and is denoted by $s(\gamma)$.
If $s(\gamma)=(a_{1},\dots,a_{n})$, then the number of perfect matchings $m(\mathcal{G}[a_{1},\dots,a_{n}])$ is called the $(k_{1},k_{2},k_{3},\sigma)$-generalized Markov (GM) length of $\gamma$, denoted by $|\gamma|$. Note that $|\gamma|$ is independent of the orientation of $\gamma$.
Example 5.4.
Set $(k_{1},k_{2},k_{3},\sigma)=(1,2,0,\mathrm{id})$. Let $\gamma$ be the red oriented generalized arc shown in Figure 16. Then
$s(\gamma)=(1,7,1,8,1,1,2,2,6,5),$
and
$|\gamma|=33848.$
Figure 16. Generalized arc $\gamma$ and sign assignments.
For any two lattice points $A=(a,b),B=(c,d)\in\mathbb{Z}^{2}$ in $\widetilde{\mathbb{R}^{2}}$, we define the $(k_{1},k_{2},k_{3},\sigma)$-generalized Markov distance. If aca-c and bdb-d are relatively prime, let $\gamma_{AB}$ denote the straight line segment from $A$ to $B$. Otherwise, $\gamma_{AB}$ is defined as a curve obtained by slightly shifting the interior of the line segment from $A$ to $B$ uniformly to the right (or left) to avoid any lattice points lying on the segment. Note that for any edge of $\widetilde{\mathbb{R}^{2}}$ whose midpoint lies on the line segment from $A$ to $B$, $\gamma_{AB}$ intersects that edge at a point shifted in the same direction as the chosen displacement. We define
$d(A,B):=\begin{cases}|\gamma_{AB}|&\text{if $A\neq B$},\\ 0&\text{if $A=B$,}\end{cases}$
which is called the $(k_{1},k_{2},k_{3},\sigma)$-generalized Markov (GM) distance between $A$ and $B$.
Remark 5.5.
Although the curve $\gamma_{AB}$ depends on whether the interior of the line segment from $A$ to $B$ is shifted to the left or right, the value of $d(A,B)$ is independent of this choice. This fact is proved by [banaian]*Lemma 8 for $k_{1}=k_{2}=k_{3}$, and we can apply the same argument to general cases.
Example 5.6.
Set $(k_{1},k_{2},k_{3},\sigma)=(1,2,0,\mathrm{id})$.
  • (1)
    Let $A=(0,0)$ and $B=(3,2)$. Then $\gamma_{AB}$ is the red curve segment shown in Figure 17. Then we have
    $s(\gamma_{AB})=(4,4,5,4)$
    and
    $d(A,B)=373.$
    Figure 17. $(1,2,0,\mathrm{id})$-GM distance between $(0,0)$ and $(3,2)$.
  • (2)
    Let $A=(0,0)$ and $B=(6,4)$. Then $\gamma_{AB}$ is the red curve segment shown in Figure 18. Then we have
    $s(\gamma_{AB})=(4,5,4,4,5,1,3,5,4,4)$
    and
    $d(A,B)=834774.$
    Note that the intersection point of the horizontal edge in $\widetilde{\mathbb{R}^{2}}$ and $\gamma_{AB}$ is slightly to the left of the midpoint of the horizontal edge, as in Example 4.4.
    Figure 18. $(1,2,0,\mathrm{id})$-GM distance between $(0,0)$ and $(6,4)$.
The distance $d(A,B)$ gives the minimal GM length among all generalized arcs connecting the lattice points $A$ and $B$, as formalized in the following theorem:
Theorem 5.7 ([banaian]*Proposition 2 and Lemma 9).
Let $\gamma$ be a generalized arc connecting lattice points $A$ and $B$ in $\widetilde{\mathbb{R}^{2}}$. Then
$|\gamma|\geq d(A,B).$
In [banaian]*Lemma 8, only the case $k_{1}=k_{2}=k_{3}$ is treated, but the theorem applies in the same way even when these three values are distinct. In the case $k_{1}=k_{2}=k_{3}=0$, Theorem 5.7 is proved by [llrs]*Theorem 3.5.
Remark 5.8.
In this paper, we use Theorem 5.7 only in the case where $\gamma_{AB}$ is a simple line segment, i.e., the differences of the first and second components of $A$ and $B$ are relatively prime.
6. Generalized Discrete Markov Spectra
In this section, we generalize Theorem 2.23 to GM numbers, and give some properties.
6.1. Main theorems and their Proofs
The main theorem of our paper is the following:
Theorem 6.1.
Fix $(k_{1},k_{2},k_{3})\in\mathbb{Z}^{3}$ and $\sigma\in\mathfrak{S}_{3}$. For any irreducible fraction $t\in[0,\infty]$, let $(n_{t},i_{t})$ denote the corresponding $(k_{1},k_{2},k_{3},\sigma)$-GM number-position pair, and let $s(t)$ be the corresponding generalized strongly admissible sequence. Then we have
$L(s(t))=\ell(^{\infty}s(t)|s(t)^{\infty})=\ell(^{\infty}s^{\ast}(\tfrac{1}{t})|s^{\ast}(\tfrac{1}{t})^{\infty})=L(s^{\ast}(\tfrac{1}{t})),$
and this value is given by
$\frac{\sqrt{((3+k_{1}+k_{2}+k_{3})n_{t}-k_{t})^{2}-4}}{n_{t}}.$
To prove Theorem 6.1, we recall the following lemma.
Lemma 6.2 (See [bombieri2]*Lemma 17).
For any finite integer sequence $S=(a_{1},\dots,a_{n})$, we have
$\ell(^{\infty}S|S^{\infty})=\frac{\sqrt{\mathrm{tr}(CF_{S})^{2}-(-1)^{n}4}}{(CF_{S})_{21}},$
where $(CF_{S})_{21}$ denotes the $(2,1)$-entry of $CF_{S}$.
Proof of Theorem 6.1.
When $t=\frac{0}{1}$ and $\frac{1}{0}$, we can check the conclusion by using Lemma 6.2. We assume that $t\in(0,\infty)$. By Theorem 4.2, Theorem 4.6, and Lemma 6.2, we obtain
$\ell(^{\infty}s(t)|s(t)^{\infty})=\frac{\sqrt{((3+k_{1}+k_{2}+k_{3})n_{t}-k_{t})^{2}-4}}{n_{t}}.$
Combining this fact with Remark 3.5 and Theorem 4.2, we deduce that
$\ell(^{\infty}s(t)|s(t)^{\infty})=\ell(^{\infty}s^{\ast}(\tfrac{1}{t})|s^{\ast}(\tfrac{1}{t})^{\infty}).$
Thus it remains to prove that
$L(s(t))=\ell(^{\infty}s(t)|s(t)^{\infty}),$
that is, $\ell(^{\infty}s(t)|s(t)^{\infty})$ is maximal among all values $\ell(P^{\ast}|Q)$ for divisions $^{\infty}s(t)^{\infty}=P^{\ast}Q$. We set $s(t)=(a_{1},a_{2},\dots,a_{n})$. The possible divisions of $^{\infty}s(t)^{\infty}$ are given by
$\dots a_{1},a_{2}\dots a_{n-1},a_{n}$ $\;|\;a_{1},a_{2},\dots,a_{n-1},a_{n}\dots,$
$\dots a_{2},a_{3}\dots a_{n},a_{1}$ $\;|\;a_{2},a_{3},\dots,a_{n},a_{1}\dots,$
$\vdots$
$\dots a_{n},a_{1}\dots a_{n-2},a_{n-1}$ $\;|\;a_{n},a_{1},\dots,a_{n-2},a_{n-1}\dots,$
therefore there are $n$ such decompositions. By Lemma 6.2, it suffices to show that $\ell(^{\infty}s(t)|s(t)^{\infty})$ is the maximum value in
$F_{t}=\left\{\frac{\sqrt{\mathrm{tr}(CF_{S})^{2}-4}}{(CF_{S})_{21}}\ \middle|\ S=(a_{k}\dots,a_{k+n-1})\text{ for some $k$ }\right\},$
where indices of $a_{i}$ are taken modulo $n$. The matrices $CF_{(a_{i},a_{i+1},\dots,a_{i-1})}$ are related by conjugation, hence all have the same trace. Therefore, the maximum value in $F_{t}$ has the minimal denominator $(CF_{(a_{i},a_{i+1},\dots,a_{i-1})})_{21}$. By (4.1), we have
$(CF_{(a_{i},a_{i+1},\dots,a_{i+n-1})})_{21}=m(\mathcal{G}[a_{i+1},\dots,a_{i+n-1}]).$
Therefore, to prove the claim, it suffices to show that $m(\mathcal{G}[a_{2},\dots,a_{n}])$ is the minimal element in
$N_{t}=\{m(\mathcal{G}[w])\mid w=(a_{k+1},a_{k+2},\dots,a_{k+n-1})\ \text{for some $k$}\ \}.$
By projecting $\widetilde{\mathbb{R}^{2}}$ to a triangulation of the once-punctured torus, $\overline{L_{t}}$ goes to a loop on the torus (see also Remark 2.12). Viewing $\overline{L_{t}}$ as a loop on this torus, for each $w=(a_{k+1},\dots,a_{k+n-1})$, let $\overline{L_{t}}(w)$ denote the segment of $\overline{L_{t}}$ producing the sign sequence $w$ by the triangle-crossing and edge-crossing rules. By modifying the endpoints of $\overline{L_{t}}(w)$ as in Table 1, we obtain an arc, which is denoted by $\widetilde{L_{t}}(w)$, such that $w$ is its sign sequence according to the triangle-crossing, edge-crossing, and endpoint rules.
Endpoint of $\overline{L_{t}}(w)$ $\widetilde{L_{t}}(w)$-Modification
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
Endpoint of $\overline{L_{t}}(w)$ $\widetilde{L_{t}}(w)$-Modification
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
Table 1. Endpoint modification
We note that the two endpoints of $\overline{L_{t}}(w)$ are connected by modification without any detour up to homotopy on the punctured torus (see Figure 19).
$\mapsto$   Refer to caption
Figure 19. Example of modification from $\overline{L_{t}}(w)$ to $\widetilde{L_{t}}(w)$
We can check this fact by enumerating all patterns of endpoint modifications. Therefore, $\widetilde{L_{t}}(w)$ can be represented as curve segment from $(0,0)$ to $(q,p)$ if $t=\frac{p}{q}$. If $\widetilde{L_{t}}(w)$ touches an edge twice in succession, first at an endpoint and then by passing through it ,or in the reverse order, then it is not a generalized arc. In this case, we modify the end of $\widetilde{L_{t}}(w)$ by deleting the initial (or final) passage and connecting it directly to the endpoint (see the figure 20).
$\mapsto$   Refer to caption
Figure 20. Example of modification from $\widetilde{L_{t}}(w)$ to $\widetilde{L_{t}}(w)_{-}$
We denote the resulting generalized arc by $\widetilde{L_{t}}(w)_{-}$. Construct the sequence associated to $\widetilde{L_{t}}(w)_{-}$ by applying the crossing rules and the endpoint rule in the same way as constructing $s(t)$. Then the resulting sequence agrees with the sequence $w=(a_{k+1},a_{k+2},\dots,a_{k+n-1})$ corresponding to $\widetilde{L_{t}}(w)$, except that its first (or last) entry is strictly smaller than $a_{k+1}$ (or $a_{k+n-1}$). We denote this sequence by
$(b_{k+1},a_{k+2},\dots,a_{k+n-2},b_{k+n-1}),$
where if $b_{k+1}=0$, then $b_{k+1}$ is removed from the beginning of the sequence; the same convention applies to $b_{k+n-1}$. In this situation,
$m(\mathcal{G}[b_{k+1},a_{k+2},\dots,a_{k+n-2},b_{k+n-1}])<m(\mathcal{G}[w])$
holds. Now we can see $\widetilde{L_{t}}(a_{2},\dots,a_{n})$ corresponds to the line segment $L_{t}$ (defined in Section 2) with endpoints $A=(0,0)$ and $B=(q,p)$. Clearly $L_{t}$ coincides with $\gamma_{AB}$. Therefore, for any $w=(a_{k+1},a_{k+2},\dots,a_{k+n-2},a_{k+n-1})$, if $\widetilde{L_{t}}(w)$ is a generalized arc, then $m(\mathcal{G}[w])$ coincides with its GM length of this arc and thus we have
$m(\mathcal{G}[a_{2},\dots,a_{n}])\leq m(\mathcal{G}[w]),$
and if $\widetilde{L_{t}}(w)$ is not a generalized arc, then $m(\mathcal{G}[b_{k+1},a_{k+2},\dots,a_{k+n-2},b_{k+n-1}])$ coincides with its GM length of $\widetilde{L_{t}}(w)_{-}$ and thus we have
$m(\mathcal{G}[a_{2},\dots,a_{n}])\leq m(\mathcal{G}[b_{k+1},a_{k+2},\dots,a_{k+n-2},b_{k+n-1}])<m(\mathcal{G}[w])$
by Theorem 5.7. Hence $m(\mathcal{G}[a_{2},\dots,a_{n}])$ attains the minimal value in $N_{t}$, which completes the proof. ∎
Therefore, we have the following corollary:
Corollary 6.3.
Fix $(k_{1},k_{2},k_{3})\in\mathbb{Z}^{3}$ and $\sigma\in\mathfrak{S}_{3}$. For any irreducible fraction $t\in[0,\infty]$, let $(n_{t},i_{t})$ denote the corresponding $(k_{1},k_{2},k_{3},\sigma)$-GM number-position pair, and let $s(t)$ be the corresponding generalized strongly admissible sequence. Then we have
$s(t)^{\infty}]\in\mathbb{Q}[\sqrt{((3+k_{1}+k_{2}+k_{3})n_{t}-k_{t})^{2}-4}]$
and
$\mathcal{L}([s(t)^{\infty}])=\mathcal{M}(Q_{s(t)})=\frac{\sqrt{((3+k_{1}+k_{2}+k_{3})n_{t}-k_{t})^{2}-4}}{n_{t}}.$
Proof.
We will prove the former statement. Let $\alpha=[s(t)^{\infty}]$. Then, we have
$\alpha=[s(t),\alpha].$
By general theory of continued fraction matrix (see [bombieri2]*Appendix A), we have
$\alpha=CF_{s(t)}(\alpha):=\dfrac{a\alpha+b}{c\alpha+d},$
where $CF_{s(t)}=\begin{bmatrix}a&b\\ c&d\end{bmatrix}$, and thus
$c\alpha^{2}-(a-d)\alpha-b=0.$
Therefore, we have
$\alpha=\dfrac{a-d\pm\sqrt{(a-d)^{2}+4bc}}{2c}=\dfrac{a-d\pm\sqrt{(a+d)^{2}-4(ad-bc)}}{2c}.$
Since $\alpha>0$, we have
$\alpha=\dfrac{a-d+\sqrt{(a+d)^{2}-4(ad-bc)}}{2c}.$
When $t=\frac{0}{1}$, we can check the conclusion directly. We assume that $t\in(0,\infty].$ By Theorem 4.2 and 4.6, we have
$\alpha=\dfrac{((3+k_{1}+k_{2}+k_{3})n_{t}-k_{t}-2u_{t}+\sqrt{((3+k_{1}+k_{2}+k_{3})n_{t}-k_{t})^{2}-4}}{2n_{t}}.$
Therefore, we have
$\alpha\in\mathbb{Q}[\sqrt{((3+k_{1}+k_{2}+k_{3})n_{t}-k_{t})^{2}-4}].$
The latter statement follows from Theorems 6.1, 2.20 and 2.22. ∎
Example 6.4.
Set $(k_{1},k_{2},k_{3},\sigma)=(1,2,0,\textrm{id})$ and $t=\tfrac{2}{5}$. Then $i_{\frac{2}{5}}=1$ and
$s\!\left(\tfrac{2}{5}\right)=(5,1,3,3,1,5,4,1,3,4).$
Thus the sequences $w$ that contribute to $m(\mathcal{G}[w])\in N_{\frac{2}{5}}$ are
$(1,3,3,1,5,4,1,3,4),\ (3,3,1,5,4,1,3,4,5),\ (3,1,5,4,1,3,4,5,1),\ (1,5,4,1,3,4,5,1,3),$
$(5,4,1,3,4,5,1,3,3),\ (4,1,3,4,5,1,3,3,1),\ (1,3,4,5,1,3,3,1,5),\ (3,4,5,1,3,3,1,5,4),$
$(4,5,1,3,3,1,5,4,1),\ (5,1,3,3,1,5,4,1,3).$
The corresponding values $m(\mathcal{G}[w])$ are
8227, 32957, 12039, 12041, 32937, 8261, 9997, 31881, 12199, 11127,8227,\ 32957,\ 12039,\ 12041,\ 32937,\ 8261,\ 9997,\ 31881,\ 12199,\ 11127,
and the minimal number in $N_{\frac{2}{5}}$ is $8227$. The corresponding arcs $\overline{L_{t}}(w)$ and $\widetilde{L_{t}}(w)$ are as in Table 2. Since
$CF_{s(\frac{2}{5})}=\begin{bmatrix}m(\mathcal{G}[5,1,3,3,1,5,4,1,3,4]&m(\mathcal{G}[5,1,3,3,1,5,4,1,3])\\ m(\mathcal{G}[1,3,3,1,5,4,1,3,4])&m(\mathcal{G}[1,3,3,1,5,4,1,3])\end{bmatrix}=\begin{bmatrix}47431&11127\\ 8227&1930\end{bmatrix}$,
we have
$s(\tfrac{2}{5})^{\infty}]=\dfrac{\sqrt{2436508317}+45501}{16454},\quad L(s(\tfrac{2}{5}))=\dfrac{\sqrt{2436508317}}{8227}.$
Moreover, we have
$Q_{s(\frac{2}{5})}=x^{2}-\frac{45501}{8227}xy-\frac{11127}{8227}y^{2},$
and thus
$\mathcal{L}\left(\frac{\sqrt{2436508317}+45501}{16454}\right)=\mathcal{M}\left(x^{2}-\frac{45501}{8227}xy-\frac{11127}{8227}y^{2}\right)=\dfrac{\sqrt{2436508317}}{8227}.$
$w$ $\overline{L_{t}}(w)$ $\widetilde{L_{t}}(w)$ $m(\mathcal{G}[w])$
$(1,3,3,1,5,4,1,3,4)$ [Uncaptioned image] [Uncaptioned image] $8227$
$(3,3,1,5,4,1,3,4,5)$ [Uncaptioned image] [Uncaptioned image] $32957$
$(3,1,5,4,1,3,4,5,1)$ [Uncaptioned image] [Uncaptioned image] $12039$
$(1,5,4,1,3,4,5,1,3)$ [Uncaptioned image] [Uncaptioned image] $12041$
$(5,4,1,3,4,5,1,3,3)$ [Uncaptioned image] [Uncaptioned image] $32937$
$(4,1,3,4,5,1,3,3,1)$ [Uncaptioned image] [Uncaptioned image] 8261
$(1,3,4,5,1,3,3,1,5)$ [Uncaptioned image] [Uncaptioned image] $9997$
$(3,4,5,1,3,3,1,5,4)$ [Uncaptioned image] [Uncaptioned image] $31881$
$(4,5,1,3,3,1,5,4,1)$ [Uncaptioned image] [Uncaptioned image] $12199$
$(5,1,3,3,1,5,4,1,3)$ [Uncaptioned image] [Uncaptioned image] $11127$
Table 2. $\overline{L_{t}}(w)$ and $\widetilde{L_{t}}(w)$
Remark 6.5.
Here are some tips for carrying out concrete example calculations:
  • (1)
    For given $(k_{1},k_{2},k_{3})\in\mathbb{Z}_{\geq 0}^{3}$ and $\sigma\in\mathfrak{S}_{3}$, The position $i_{t}$ corresponds to the edge of $\widetilde{\mathbb{R}^{n}}$ through which $L_{t}$ passes exactly at its midpoint. Namely, if this edge is horizontal, then $i_{t}=\sigma(1)$; if it is diagonal, then $i_{t}=\sigma(2)$; and if it is vertical, then $i_{t}=\sigma(3)$. Moreover, $k_{t}$ coincides with the number of signs associated with this edge.
  • (2)
    For an integer sequence $w=(b_{1},\dots,b_{n})$, $m(\mathcal{G}[w])$ corresponds to the denominator of $b_{1},\dots,b_{n}]$ and the numerator of $b_{2},\dots,b_{n}]$ by Theorem 4.11.
6.2. Properties of Generalized Discrete Markov Spectra
For any $n\in\mathbb{Z}_{\geq 1}$ and $i\in\{1,2,3\}$, we define
$\Delta(n,i):=((3+k_{1}+k_{2}+k_{3})n-k_{i})^{2}-4.$
We denote
$\mathcal{M}_{k_{1},k_{2},k_{3},\sigma}:=\left\{\frac{\sqrt{\Delta(n_{t},i_{t})}}{n_{t}}\ \middle|\ \begin{aligned} (n_{t},i_{t})\text{ is a $(k_{1},k_{2},k_{3},\sigma)$-GM number-position pair}\end{aligned}\right\}$
and consider
$\mathcal{M}_{k_{1},k_{2},k_{3}}:=\bigcup_{\sigma\in\mathfrak{S}_{3}}\mathcal{M}_{k_{1},k_{2},k_{3},\sigma}.$
The set $\mathcal{M}_{k_{1},k_{2},k_{3}}$ is called the $(k_{1},k_{2},k_{3})$-generalized discrete Markov spectrum. We note that for any $\sigma\in\mathfrak{S}_{3}$, we have $\mathcal{M}_{k_{1},k_{2},k_{3}}=\mathcal{M}_{k_{\sigma(1)},k_{\sigma(2)},k_{\sigma(3)}}.$ Moreover, if $k_{1}=k_{2}=k_{3}=0$, then $\mathcal{M}_{0,0,0}$ coincides with the ordinary discrete Markov spectrum $\mathcal{M}_{d}$.
By Corollary 6.3, we have the following corollary:
Corollary 6.6.
For any $k_{1},k_{2},k_{3}\in\mathbb{Z}_{\geq 0}$, $\mathcal{M}_{k_{1},k_{2},k_{3}}\subset\mathcal{L}$ and $\mathcal{M}_{k_{1},k_{2},k_{3}}\subset\mathcal{M}$ hold.
By the duality $(n_{t},i_{t})=(n^{\ast}_{\frac{1}{t}},i^{\ast}_{\frac{1}{t}})$, even if $\sigma$ is restricted to the alternating group $\mathfrak{A}_{3}$, the set $\mathcal{M}_{k_{1},k_{2},k_{3}}$ remains unchanged. Furthermore, even if we do not impose any restriction on $\sigma$ but instead restrict $t$ to the interval $[0,1]$, the set $\mathcal{M}_{k_{1},k_{2},k_{3}}$ remains unchanged. Moreover, if two among $k_{1},k_{2},k_{3}$ are equal, then restricting $\sigma$ to $\mathfrak{A}_{3}$ and $t$ to $[0,1]$ does not alter $\mathcal{M}_{k_{1},k_{2},k_{3}}$.
For any $(k_{1},k_{2},k_{3})\in\mathbb{Z}_{\geq 0}^{3}$, the minimal $(k_{1},k_{2},k_{3})$-number is $1$. Therefore,
$\mathcal{M}_{k_{1},k_{2},k_{3}}\subset[2+k_{i}+k_{j},3+k_{1}+k_{2}+k_{3}],$
where $k_{i},k_{j}$ are the two smallest elements among $\{k_{1},k_{2},k_{3}\}$.
Example 6.7.
We list some quadratic irrational numbers $\alpha$ and the corresponding values $\mathcal{L}(\alpha)$ obtained from Corollary 6.3. Table 3 shows the case $(k_{1},k_{2},k_{3})=(0,0,0)$, Tables 4, 5, and 6 show the case $(k_{1},k_{2},k_{3})=(0,0,1)$, Tables 7, 8 and 9 shows the case $(k_{1},k_{2},k_{3})=(0,1,1)$. Table 10 shows the case $(k_{1},k_{2},k_{3})=(1,1,1)$, and Table 11 shows the case $(k_{1},k_{2},k_{3})=(2,2,2)$. Table 12 show the case $(k_{1},k_{2},k_{3},\sigma)=(1,2,0,\mathrm{id})$. For each list, the elements are arranged in ascending order of the corresponding GM numbers.
$t$ $s(t)$ $\alpha=[s(t)^{\infty}]$ $n_{t}$ $\mathcal{L}(\alpha)$
$\dfrac{0}{1}$ $(1,1)$ $\dfrac{\sqrt{5}+1}{2}$ 11 $\sqrt{5}$
$\dfrac{1}{1}$ $(2,2)$ $\sqrt{2}+1$} 22 $2\sqrt{2}$
$\dfrac{1}{2}$ $(2,1,1,2)$ $\dfrac{\sqrt{221}+11}{10}$ 55 $\dfrac{\sqrt{221}}{5}$
$\dfrac{1}{3}$ $(2,1,1,1,1,2)$ $\dfrac{\sqrt{1517}+29}{26}$ $13$ $\dfrac{\sqrt{1517}}{13}$
$\dfrac{2}{3}$ $(2,1,1,2,2,2)$ $\dfrac{\sqrt{7565}+63}{58}$ $29$ $\dfrac{\sqrt{7565}}{29}$
$\dfrac{1}{4}$ $(2,1,1,1,1,1,1,2)$ $\dfrac{5\sqrt{26}+19}{17}$ $34$ $\dfrac{10\sqrt{26}}{17}$
$\dfrac{1}{5}$ $(2,1,1,1,1,1,1,1,1,2)$ $\dfrac{\sqrt{71285}+199}{178}$ $89$ $\dfrac{\sqrt{71285}}{89}$
$\dfrac{3}{4}$ $(2,1,1,2,2,2,2,2)$ $\dfrac{\sqrt{257045}+367}{338}$ $169$ $\dfrac{\sqrt{257045}}{169}$
Table 3. Case $(k_{1},k_{2},k_{3})=(0,0,0)$
$t$ $s(t)$ $\alpha=[s(t)^{\infty}]$ $n_{t}$ $\mathcal{L}(\alpha)$
$\dfrac{0}{1}$ $(2,1)$ $\sqrt{3}+1$ 11 $2\sqrt{3}$
$\dfrac{1}{1}$ $(3,2)$ $\dfrac{\sqrt{15}+3}{2}$ 22 $\sqrt{15}$
$\dfrac{1}{2}$ $(3,1,1,3)$ $\dfrac{5\sqrt{29}+23}{14}$ 77 $\dfrac{5\sqrt{29}}{7}$
$\dfrac{1}{3}$ $(3,1,2,1,1,3)$ $\dfrac{7\sqrt{51}+43}{25}$ $25$ $\dfrac{14\sqrt{51}}{25}$
$\dfrac{2}{3}$ $(3,1,1,3,3,2)$ $\dfrac{\sqrt{11235}+83}{53}$ $53$ $\dfrac{2\sqrt{11235}}{53}$
$\dfrac{1}{4}$ $(3,1,2,1,1,2,1,3)$ $\dfrac{\sqrt{15293}+107}{62}$ $93$ $\dfrac{\sqrt{15293}}{31}$
$\dfrac{1}{5}$ $(3,1,2,1,2,1,1,2,1,3)$ $\dfrac{3\sqrt{53207}+599}{346}$ $346$ $\dfrac{3\sqrt{53207}}{346}$
$\dfrac{3}{4}$ $(3,1,1,3,2,3,3,2)$ $\dfrac{\sqrt{308765}+435}{278}$ $417$ $\dfrac{2\sqrt{308765}}{278}$
Table 4. Case $(k_{1},k_{2},k_{3},\sigma)=(0,0,1,\mathrm{id})$
$t$ $s(t)$ $\alpha=[s(t)^{\infty}]$ $n_{t}$ $\mathcal{L}(\alpha)$
$\dfrac{0}{1}$ $(2,1)$ $\sqrt{3}+1$ 11 $2\sqrt{3}$
$\dfrac{1}{1}$ $(3,3)$ $\dfrac{\sqrt{13}+3}{2}$ 3 $\sqrt{13}$
$\dfrac{1}{2}$ $(3,1,2,3)$ $\dfrac{\sqrt{399}+17}{10}$ $10$ $\dfrac{\sqrt{399}}{5}$
$\dfrac{1}{3}$ $(3,1,2,2,1,3)$ $\dfrac{\sqrt{21605}+127}{74}$ $37$ $\dfrac{\sqrt{21605}}{37}$
$\dfrac{2}{3}$ $(3,1,2,3,3,3)$ $\dfrac{\sqrt{47523}+185}{109}$ $109$ $\dfrac{2\sqrt{47523}}{109}$
$\dfrac{1}{4}$ $(3,1,2,1,2,2,1,3)$ $\dfrac{5\sqrt{3003}+237}{137}$ $137$ $\dfrac{10\sqrt{3003}}{137}$
$\dfrac{1}{5}$ $(3,1,2,1,2,2,1,2,1,3)$ $\dfrac{\sqrt{4173845}+1769}{1022}$ $511$ $\dfrac{\sqrt{4173845}}{511}$
$\dfrac{3}{4}$ $(3,1,2,3,3,3,3,3)$ $\dfrac{\sqrt{5654883}+2018}{1189}$ $1189$ $\dfrac{2\sqrt{5654883}}{1189}$
Table 5. Case $(k_{1},k_{2},k_{3},\sigma)=(0,0,1,(1\ 2\ 3))$
$t$ $s(t)$ $\alpha=[s(t)^{\infty}]$ $n_{t}$ $\mathcal{L}(\alpha)$
$\dfrac{0}{1}$ $(1,1)$ $\dfrac{\sqrt{5}+1}{2}$ 11 $\sqrt{5}$
$\dfrac{1}{1}$ $(3,2)$ $\dfrac{\sqrt{15}+3}{2}$ 22 $\sqrt{15}$
$\dfrac{1}{2}$ $(3,1,1,2)$ $\dfrac{3\sqrt{11}+8}{5}$ 55 $\dfrac{6\sqrt{11}}{5}$
$\dfrac{1}{3}$ $(3,1,1,1,1,2)$ $\dfrac{15\sqrt{3}+21}{13}$ $13$ $\dfrac{30\sqrt{3}}{13}$
$\dfrac{1}{4}$ $(3,1,1,1,1,1,1,2)$ $\dfrac{\sqrt{4623}+55}{34}$ $34$ $\dfrac{\sqrt{4623}}{17}$
$\dfrac{2}{3}$ $(3,1,1,3,2,2)$ $\dfrac{\sqrt{2669}+41}{26}$ $39$ $\dfrac{\sqrt{2669}}{13}$
$\dfrac{1}{5}$ $(3,1,1,1,1,1,1,1,1,2)$ $\dfrac{\sqrt{31683}+144}{89}$ $89$ $\dfrac{2\sqrt{31683}}{89}$
$\dfrac{1}{6}$ $(3,1,1,1,1,1,1,1,1,1,1,2)$ $\dfrac{\sqrt{217155}+377}{233}$ $233$ $\dfrac{2\sqrt{217155}}{233}$
Table 6. Case $(k_{1},k_{2},k_{3},\sigma)=(0,0,1,(1\ 3\ 2))$
$t$ $s(t)$ $\alpha=[s(t)^{\infty}]$ $n_{t}$ $\mathcal{L}(\alpha)$
$\dfrac{0}{1}$ $(3,1)$ $\dfrac{\sqrt{21}+3}{2}$ 11 $\sqrt{21}$
$\dfrac{1}{1}$ $(4,3)$ $\dfrac{4\sqrt{3}+6}{3}$ 33 $\dfrac{8\sqrt{3}}{3}$
$\dfrac{1}{2}$ $(4,1,2,4)$ $\dfrac{\sqrt{1023}+29}{13}$ $13$ $\dfrac{2\sqrt{1023}}{13}$
$\dfrac{1}{3}$ $(4,1,3,2,1,4)$ $\dfrac{3\sqrt{2567}+139}{61}$ $61$ $\dfrac{6\sqrt{2567}}{61}$
$\dfrac{2}{3}$ $(4,1,2,4,4,3)$ $\dfrac{\sqrt{49506}+195}{89}$ $178$ $\dfrac{2\sqrt{49506}}{89}$
$\dfrac{1}{4}$ $(4,1,3,1,2,3,1,4)$ $\dfrac{44\sqrt{273}+666}{291}$ $291$ $\dfrac{88\sqrt{273}}{291}$
$\dfrac{1}{5}$ $(4,1,3,1,3,2,1,3,1,4)$ $\dfrac{531\sqrt{43}+3191}{1393}$ $1393$ $\dfrac{1062\sqrt{43}}{1393}$
$\dfrac{3}{4}$ $(4,1,2,4,3,4,4,3)$ $\dfrac{2\sqrt{9600702}+5431}{2479}$ $2479$ $\dfrac{4\sqrt{9600702}}{2479}$
Table 7. Case $(k_{1},k_{2},k_{3},\sigma)=(0,1,1,\mathrm{id})$
$t$ $s(t)$ $\alpha=[s(t)^{\infty}]$ $n_{t}$ $\mathcal{L}(\alpha)$
$\dfrac{0}{1}$ $(2,1)$ $\sqrt{3}+1$ 11 $2\sqrt{3}$
$\dfrac{1}{1}$ $(4,3)$ $\dfrac{4\sqrt{3}+6}{3}$ 33 $\dfrac{8\sqrt{3}}{3}$
$\dfrac{1}{2}$ $(4,1,2,3)$ $\dfrac{2\sqrt{39}+11}{5}$ $10$ $\dfrac{4\sqrt{39}}{5}$
$\dfrac{1}{3}$ $(4,1,2,2,1,3)$ $\dfrac{\sqrt{8463}+82}{37}$ $37$ $\dfrac{2\sqrt{8463}}{37}$
$\dfrac{1}{4}$ $(4,1,2,1,2,2,1,3)$ $\dfrac{\sqrt{469221}+611}{274}$ $137$ $\dfrac{\sqrt{469221}}{137}$
$\dfrac{2}{3}$ $(4,1,2,4,3,3)$ $\dfrac{2\sqrt{30102}+305}{139}$ $139$ $\dfrac{4\sqrt{30102}}{139}$
$\dfrac{1}{5}$ $(4,1,2,1,2,2,1,2,1,3)$ $\dfrac{6\sqrt{45298}+1140}{511}$ $511$ $\dfrac{12\sqrt{45298}}{511}$
$\dfrac{2}{5}$ $(4,1,2,2,1,4,3,1,2,3)$ $\dfrac{22\sqrt{43662}+4050}{1839}$ $1839$ $\dfrac{44\sqrt{43662}}{1839}$
Table 8. Case $(k_{1},k_{2},k_{3},\sigma)=(0,1,1,(1\ 2\ 3))$
$t$ $s(t)$ $\alpha=[s(t)^{\infty}]$ $n_{t}$ $\mathcal{L}(\alpha)$
$\dfrac{0}{1}$ $(2,1)$ $\sqrt{3}+1$ 11 $2\sqrt{3}$
$\dfrac{1}{1}$ $(4,2)$ $\sqrt{6}+2$ 22 $2\sqrt{6}$
$\dfrac{1}{2}$ $(4,1,1,3)$ $\dfrac{12\sqrt{2}+15}{7}$ 77 $\dfrac{24\sqrt{2}}{7}$
$\dfrac{1}{3}$ $(4,1,2,1,1,3)$ $\dfrac{\sqrt{15621}+111}{50}$ $25$ $\dfrac{\sqrt{15621}}{25}$
$\dfrac{2}{3}$ $(4,1,1,4,3,2)$ $\dfrac{4\sqrt{1743}+138}{67}$ $67$ $\dfrac{8\sqrt{1743}}{67}$
$\dfrac{1}{4}$ $(4,1,2,1,1,2,1,3)$ $\dfrac{\sqrt{53823}+207}{93}$ $93$ $\dfrac{2\sqrt{53823}}{93}$
$\dfrac{1}{5}$ $(4,1,2,1,2,1,1,2,1,3)$ $\dfrac{12\sqrt{1299}+386}{173}$ $346$ $\dfrac{24\sqrt{1299}}{173}$
$\dfrac{3}{4}$ $(4,1,1,4,2,3,4,2)$ $\dfrac{\sqrt{2729103}+1356}{661}$ $661$ $\dfrac{2\sqrt{2729103}}{661}$
Table 9. Case $(k_{1},k_{2},k_{3},\sigma)=(0,1,1,(1\ 3\ 2))$
$t$ $s(t)$ $\alpha=[s(t)^{\infty}]$ $n_{t}$ $\mathcal{L}(\alpha)$
$\dfrac{0}{1}$ $(3,1)$ $\dfrac{\sqrt{21}+3}{2}$ 11 $\sqrt{21}$
$\dfrac{1}{1}$ $(5,3)$ $\dfrac{\sqrt{285}+15}{6}$ 33 $\dfrac{\sqrt{285}}{3}$
$\dfrac{1}{2}$ $(5,1,2,4)$ $\dfrac{5\sqrt{237}+71}{26}$ $13$ $\dfrac{5\sqrt{237}}{13}$
$\dfrac{1}{3}$ $(5,1,3,2,1,4)$ $\dfrac{11\sqrt{1101}+339}{122}$ $61$ $\dfrac{11\sqrt{1101}}{61}$
$\dfrac{2}{3}$ $(5,1,2,5,4,3)$ $\dfrac{\sqrt{1692597}+1167}{434}$ $217$ $\dfrac{\sqrt{1692597}}{217}$
$\dfrac{1}{4}$ $(5,1,3,1,2,3,1,4)$ $\dfrac{\sqrt{3045021}+1623}{582}$ $291$ $\dfrac{\sqrt{3045021}}{291}$
$\dfrac{1}{5}$ $(5,1,3,1,3,2,1,3,1,4)$ $\dfrac{\sqrt{69839445}+7775}{2786}$ $1393$ $\dfrac{\sqrt{69839445}}{1393}$
$\dfrac{3}{4}$ $(5,1,2,5,3,4,5,3)$ $\dfrac{\sqrt{485629365}+19735}{7346}$ $3673$ $\dfrac{\sqrt{485629365}}{3673}$
Table 10. Case $(k_{1},k_{2},k_{3})=(1,1,1)$
$t$ $s(t)$ $\alpha=[s(t)^{\infty}]$ $n_{t}$ $\mathcal{L}(\alpha)$
$\dfrac{0}{1}$ $(5,1)$ $\dfrac{3\sqrt{5}+5}{2}$ 11 $3\sqrt{5}$
$\dfrac{1}{1}$ $(8,4)$ $\dfrac{{3\sqrt{2}+4}}{2}$ 44 $6\sqrt{2}$
$\dfrac{1}{2}$ $(8,1,3,6)$ $\dfrac{3\sqrt{221}+43}{10}$ $25$ $\dfrac{3\sqrt{221}}{5}$
$\dfrac{1}{3}$ $(8,1,5,3,1,6)$ $\dfrac{3\sqrt{1517}+113}{26}$ $169$ $\dfrac{3\sqrt{1517}}{13}$
$\dfrac{2}{3}$ $(8,1,3,8,6,4)$ $\dfrac{3\sqrt{7565}+247}{58}$ $841$ $\dfrac{3\sqrt{7565}}{29}$
$\dfrac{1}{4}$ $(8,1,5,1,3,5,1,6)$ $\dfrac{15\sqrt{26}+74}{17}$ $1156$ $\dfrac{30\sqrt{26}}{17}$
$\dfrac{1}{5}$ $(8,1,5,1,5,3,1,5,1,6)$ $\dfrac{3\sqrt{71285}+775}{178}$ $7921$ $\dfrac{3\sqrt{71285}}{89}$
$\dfrac{3}{4}$ $(8,1,3,8,4,6,8,4)$ $\dfrac{3\sqrt{257045}+1439}{338}$ $28561$ $\dfrac{3\sqrt{257045}}{169}$
Table 11. Case $(k_{1},k_{2},k_{3})=(2,2,2)$
$t$ $s(t)$ $\alpha=[s(t)^{\infty}]$ $n_{t}$ $\mathcal{L}(\alpha)$
$\dfrac{0}{1}$ $(3,1)$ $\dfrac{\sqrt{2}1+3}{2}$ 11 $\sqrt{21}$
$\dfrac{1}{1}$ $(5,4)$ $\dfrac{{\sqrt{30}+5}}{2}$ 44 $\sqrt{30}$
$\dfrac{1}{2}$ $(5,1,3,4)$ $\dfrac{10\sqrt{26}+47}{17}$ $17$ $\dfrac{20\sqrt{26}}{17}$
$\dfrac{1}{3}$ $(5,1,3,3,1,4)$ $\dfrac{\sqrt{723}+25}{9}$ $81$ $\dfrac{2\sqrt{723}}{9}$
$\dfrac{2}{3}$ $(5,1,3,5,4,4)$ $\dfrac{\sqrt{5004165}+2061}{746}$ $373$ $\dfrac{\sqrt{5004165}}{373}$
$\dfrac{1}{4}$ $(5,1,3,1,3,3,1,4)$ $\dfrac{\sqrt{1340963}+1077}{386}$ $386$ $\dfrac{\sqrt{1340963}}{193}$
$\dfrac{1}{5}$ $(5,1,3,1,3,3,1,5,1,4)$ $\dfrac{\sqrt{16635}+120}{43}$ $1849$ $\dfrac{2\sqrt{16635}}{43}$
$\dfrac{3}{4}$ $(5,1,3,5,4,4,5,4)$ $\dfrac{2\sqrt{150737006}+22602}{8185}$ $8185$ $\dfrac{4\sqrt{150737006}}{8185}$
Table 12. Case $(k_{1},k_{2},k_{3},\sigma)=(1,2,0,\mathrm{id})$
We now describe a special relation between $\mathcal{M}_{0,0,0}$ and $\mathcal{M}_{2,2,2}$. The following relation is a known fact:
Proposition 6.8 ([gyomatsu]*Theorem 11).
If $(a,b,c)$ is a $(0,0,0)$-GM triple, then $(a^{2},b^{2},c^{2})$ is a $(2,2,2)$-GM triple. Conversely, if $(A,B,C)$ is a $(2,2,2)$-GM triple, then $(\sqrt{A},\sqrt{B},\sqrt{C})$ is a $(0,0,0)$-GM triple.
Using Proposition 6.8, we have the following:
Theorem 6.9.
If $r\in\mathcal{M}_{0,0,0}$, then $3r\in\mathcal{M}_{2,2,2}$. Conversely, if $R\in\mathcal{M}_{2,2,2}$, then $\frac{R}{3}\in\mathcal{M}_{0,0,0}$.
Proof.
We first prove the forward implication. Since $r\in\mathcal{M}_{0,0,0}$, there exists a $(0,0,0)$-GM number $n$ such that
$r=\frac{\sqrt{9n^{2}-4}}{n}.$
Thus
$3r=\frac{3\sqrt{9n^{2}-4}}{n}=\frac{\sqrt{81n^{4}-36n^{2}}}{n^{2}}=\frac{\sqrt{(9n^{2}-2)^{2}-4}}{n^{2}}.$
By Proposition 6.8, $n^{2}$ is a $(2,2,2)$-GM number, and hence $3r\in\mathcal{M}_{2,2,2}$. The reverse implication follows by reversing the above computation. ∎
Next, we will give a characterization of generalized discrete Markov spectra in the transition interval:
Theorem 6.10.
Set
$\mathcal{M}^{\prime}=\bigcup_{(k_{1},k_{2},k_{3})\in\mathbb{Z}^{3}_{\geq 0}}\mathcal{M}_{k_{1},k_{2},k_{3}}.$
Then we have
$\mathcal{M}^{\prime}\cap[3,c_{F})=(\mathcal{M}_{0,0,1}\setminus\{\sqrt{5}\})\cup\{2\sqrt{5}\},$
where $c_{F}$ is the Freiman constant i.e.
$c_{F}:=\dfrac{2221564096+283748\sqrt{462}}{491993569}\approx 4.5278295661\dots.$
Proof.
Without loss of generality, we assume that $k_{1}\leq k_{2}\leq k_{3}$. First, we will prove $\mathcal{M}^{\prime}\cap[3,c_{F})\supset(\mathcal{M}_{0,0,1}\setminus\{\sqrt{5}\})\cup\{2\sqrt{5}\}$. Clearly, $\mathcal{M}_{0,0,1}\setminus\{\sqrt{5}\}\subset\mathcal{M}^{\prime}$ holds. Moreover, we have $2\sqrt{5}\in\mathcal{M}_{0,0,2}\subset\mathcal{M}^{\prime}$. Indeed, $(4,3)$ is a $(0,0,2,\mathrm{(1\ 2\ 3)})$-GM number-position pair, and the corresponding value in $\mathcal{M}_{0,0,2}$ is
$\frac{\sqrt{(5\cdot 4-2)^{2}-4}}{4}=\frac{\sqrt{320}}{4}=2\sqrt{5}.$
We will prove $(\mathcal{M}_{0,0,1}\setminus\{\sqrt{5}\})\cup\{2\sqrt{5}\}\subset[3,c_{F})$. Clearly, $2\sqrt{5}\in[3,c_{F})$. The functions
$f(x)=\frac{\sqrt{(4x)^{2}-4}}{x},\quad g(x)=\frac{\sqrt{(4x-1)^{2}-4}}{x}$
are both increasing monotonically on $x\geq 1$, and
$\lim_{x\to\infty}f(x)=\lim_{x\to\infty}g(x)=4.$
Therefore, we have $\mathcal{M}_{0,0,1}\subset[0,c_{F}]$. The minimal value in $\mathcal{M}_{0,0,1}\setminus\{\sqrt{5}\}$ is $2\sqrt{3}$ by Tables 4, 5, 6, therefore we have $\mathcal{M}_{0,0,1}\setminus\{\sqrt{5}\}\subset[3,c_{F})$.
Next, we will prove $\mathcal{M}^{\prime}\cap[3,c_{F})\subset\mathcal{(}\mathcal{M}_{0,0,1}\setminus\{\sqrt{5}\})\cup\{2\sqrt{5}\}$. Clearly,
$\mathcal{M}_{0,0,0}\cap[3,c_{F})=\emptyset,\text{ and }\mathcal{M}_{0,0,1}\cap[3,c_{F})=\mathcal{M}_{0,0,1}\setminus{\{\sqrt{5}\}}.$
We assume $k_{1}=k_{2}=0$ and $k_{3}=k\in\mathbb{Z}_{\geq 2}$. We will determine the order of small values in $\mathcal{M}_{0,0,k}$. In this case, all $(0,0,k)$-GM number-position pairs are in
$\left\{(n_{t},i_{t})\in v\ \middle|\ v\in\bigcup_{\sigma\in\mathfrak{A}_{3}}\mathrm{M}\mathbb{T}(0,0,k,\sigma),t\in[0,1]\right\}.$
First, we assume that $k_{t}=0$ for $t\in[0,1]$. Then, the minimum value and the second minimum value is as follows:
$\sigma$ Minimum $n_{t}$ Second minimum $n_{t}$
$\mathrm{id}$ $n_{\frac{0}{1}}=1$ $n_{\frac{1}{1}}=2$
(1 2 3)(1\ 2\ 3) $n_{\frac{0}{1}}=1$ $n_{\frac{1}{2}}=k^{2}+4k+5$
(1 3 2)(1\ 3\ 2) $n_{\frac{1}{1}}=2$ $n_{\frac{1}{2}}=5$
Hence, in this range, the $(0,0,k)$-GM numbers are arranged in ascending order as $1,2,5$. The corresponding elements of $\mathcal{M}_{0,0,k}$ to $n=1,2,5$ are
$\sqrt{k^{2}+6k+5},\sqrt{k^{2}+6k+8},\sqrt{k^{2}+6k+\dfrac{221}{25}}$
Next, we assume that $k_{t}=k$ for $t\in[0,1]$. Then, the minimum value and the second minimum value is as follows:
$\sigma$ Minimum $n_{t}$ Second minimum $n_{t}$
$\mathrm{id}$ $n_{\frac{1}{2}}=2k+5$ $n_{\frac{1}{4}}=2k^{3}+16k^{2}+41k+34$
(1 2 3)(1\ 2\ 3) $n_{\frac{1}{1}}=k+2$ $n_{\frac{1}{3}}=2k^{2}+10k+13$
(1 3 2)(1\ 3\ 2) $n_{\frac{0}{1}}=1$ $n_{\frac{2}{3}}=10k+29$
Hence, in this range, the $(0,0,k)$-GM numbers are arranged in ascending order as $1,k+2,2k+5$. The corresponding elements of $\mathcal{M}_{0,0,k}$ to $n=1,k+2,2k+5$ are
$\sqrt{5},\sqrt{k^{2}+4k+8},\sqrt{k^{2}+5k+\dfrac{35}{4}+\dfrac{9}{4(2k+5)^{2}}}.$
Therefore, first two minimum values in $\mathcal{M}_{0,0,k}$ are
$\sqrt{5},\sqrt{k^{2}+4k+8}.$
if $k\geq 2$. The third minimum value is
$\begin{cases}\sqrt{k^{2}+6k+5}\quad&\text{if }k=2,3,\vskip 8.53581pt\\ \sqrt{k^{2}+5k+\dfrac{35}{4}+\dfrac{9}{4(2k+5)^{2}}}\quad&\text{otherwise.}\end{cases}$
Among these values, the only one lying in $3,c_{F})$ is $\sqrt{k^{2}+4k+8}=2\sqrt{5}$, which occurs when $k=2$. Therefore, we have
$\mathcal{M}_{0,0,2}\cap[3,c_{F})=\{2\sqrt{5}\},$
and for any $k\geq 3$,
$\mathcal{M}_{0,0,k}\cap[3,c_{F})=\emptyset.$
Next, we assume that $k_{1}=0$ and $k_{2}=k_{3}=1$. By Tables 7, 8, 9, we have
$\mathcal{M}_{0,1,1}\cap[3,c_{F})=\{2\sqrt{3}\},$
and $2\sqrt{3}\in\mathcal{M}_{0,0,1}$.
Let us prove the remaining cases. We set $K:=3+k_{1}+k_{2}+k_{3}$. The remaining cases all satisfy $K\geq 6$ and the condition that $k_{i}\leq K-4$ for every $i$. Hence, from now on we shall assume these conditions. If the following inequality holds
$\dfrac{\sqrt{(Kn_{t}-k_{t})^{2}-4}}{n_{t}}<5,$
then by assumption we have
$\dfrac{\sqrt{(Kn_{t}-K+4)^{2}-4}}{n_{t}}\leq\dfrac{\sqrt{(Kn_{t}-k_{t})^{2}-4}}{n_{t}}<5.$
From the two sides of this inequality we obtain the following inequality:
$(K^{2}-25)n_{t}^{2}+(-2K^{2}+8K)n_{t}+(K^{2}-8K+12)<0.$
Since $K\geq 6$, we have $K^{2}-25>0$ and thus
$\dfrac{K(K-4)-\sqrt{29K^{2}-200K+300}}{K^{2}-25}<n_{t}<\dfrac{K(K-4)+\sqrt{29K^{2}-200K+300}}{K^{2}-25}.$
Now, the function $F(x)=\dfrac{x(x-4)+\sqrt{29x^{2}-200x+300}}{x^{2}-25}$ is monotonically decreasing on $x\geq 6$ (check the region satisfying $F^{\prime}(x)<0$). Therefore, we have
$n_{t}<\dfrac{6(6-4)+\sqrt{29\cdot 36-200\cdot 6+300}}{36-25}=\frac{8}{3}.$
Therefore, if
$\dfrac{\sqrt{((3+k_{1}+k_{2}+k_{3})n_{t}-k_{t})^{2}-4}}{n_{t}}\in[3,c_{F}),$
then $n_{t}$ must be $1$ or $2$. First, we assume $n_{t}=1$. If $\sqrt{(3+k_{1}+k_{2}+k_{3}-k_{t})^{2}-4}\in[3,c_{F})$, then $k_{1}+k_{2}+k_{3}-k_{t}=1$. We have thus $k_{1}=0$, $k_{2}=1$, $k_{t}=k_{3}$ and
$\sqrt{(3+k_{1}+k_{2}+k_{3}-k_{t})^{2}-4}=2\sqrt{3}\in\mathcal{M}_{0,0,1}.$
Second, we assume that $n_{t}=2$. Now, $k_{t}$ must be 0 and thus $k_{1}=0$. Then we have
$\dfrac{\sqrt{((3+k_{1}+k_{2}+k_{3})n_{t}-k_{i})^{2}-4}}{n_{t}}=\dfrac{\sqrt{(2\cdot(3+k_{2}+k_{3}))^{2}-4}}{2}$
From the assumption, $k_{2}+k_{3}$ must be at least $3$, but then
$\dfrac{\sqrt{(2\cdot(3+k_{2}+k_{3}))^{2}-4}}{2}\notin[3,c_{F}).$
By the above argument, if $k\geq 2$, we have
$\mathcal{M}_{0,1,k}\cap[3,c_{F})=\{2\sqrt{3}\}\subset\mathcal{M}_{0,0,1},$
and if $(k_{1},k_{2},k_{3})$ does not have the form $(0,0,0),(0,0,k)$ or $(0,1,k)$, then we have
$\mathcal{M}_{k_{1},k_{2},k_{3}}\cap[3,c_{F})=\emptyset.$
This finishes the proof. ∎
Remark 6.11.
The set $\mathcal{M}^{\prime}\cap[3,c_{F})$ does not cover all the elements of the Markov-Lagrange spectrum contained in the transition interval. In other words, the strict inclusion
$\mathcal{M}^{\prime}\cap[3,c_{F})\subsetneq\mathcal{L}\cap[3,c_{F})$
holds. We will show this fact. We set $s=(1,1,1,2,2,2)$, and
$\alpha=[(1,1,1,2,2,2)^{\infty}]=\frac{2\sqrt{210}+17}{29}.$
Using Theorem 2.20 and Lemma 6.2, we have
$\mathcal{L}(\alpha)=L(1,1,1,2,2,2)=\frac{4\sqrt{210}}{19}\approx 3.05081615709$...,
and $\mathcal{L}(\alpha)\in\mathcal{L}\cap[3,c_{F})$. On the other hand, the minimal value in $\mathcal{M}^{\prime}\cap[3,c_{F})$ is
$2\sqrt{3}\approx 3.46410161514\dots,$
and $\mathcal{L}(\alpha)\notin\mathcal{M}^{\prime}\cap[3,c_{F})$.
Theorem 2.25 states that the discrete Markov spectrum coincides with the part of the Markov spectrum less than $3$. However, a corresponding theorem for the GM numbers has not yet been discovered.
Question 6.12.
Does there exist a characterization of $\mathcal{M}_{k_{1},k_{2},k_{3}}$ (or the union set of them) that generalizes Markov’s theorem?
7. Formulation of Generalized Uniqueness Conjecture
In this section, building on the results obtained in the previous section, we discuss a natural generalization of Frobenius’s uniqueness conjecture. The original conjecture is stated as follows.
Conjecture 7.1 ([frobenius]).
For any Markov number $c$, there exists a unique Markov triple $(a,b,c)$ such that $a\leq b\leq c$.
This conjecture is equivalent to the following (see [aig]*Corollary 9.30):
Conjecture 7.2.
For any element $L\in\mathcal{M}_{0,0,0}$, if $L=\mathcal{L}(\alpha)=\mathcal{L}(\beta)$, $\alpha$ and $\beta$ are $GL(2,\mathbb{Z})$-equivalent. That is, there exists
$\begin{bmatrix}a&b\\ c&d\end{bmatrix}\in GL(2,\mathbb{Z})\quad\text{such that}\quad\alpha=\frac{a\beta+b}{c\beta+d}.$
As a generalization of Conjecture 7.1, the author and Matsushita [gyomatsu] posed the following question:
Question 7.3.
For any $(k_{1},k_{2},k_{3})$-GM number $c$, is there a unique $(k_{1},k_{2},k_{3})$-GM triple $(a,b,c)$ such that $a\leq b\leq c$?
One may also consider the following question as a generalization of Conjecture 7.2.
Question 7.4.
For any element $L\in\mathcal{M}_{k_{1},k_{2},k_{3}}$, if $L=\mathcal{L}(\alpha)=\mathcal{L}(\beta)$, are $\alpha$ and $\beta$ $GL(2,\mathbb{Z})$-equivalent?
Note that Question 7.3 and Question 7.4 are not equivalent statements.
A counterexample of Question 7.3 arises when $k_{1},k_{2},k_{3}$ are distinct ([gyomatsu]*Remark 10). In fact, both $(1,81,17)$ and $(7,81,2)$ are $(1,2,0)$-GM triples (and moreover positive integer solutions of the $(1,2,0)$-GM equation).
According to [hurwitz], Question 7.4 has the negative answer for $\mathcal{M}\setminus\mathcal{M}_{0,0,0}$. Here we shall introduce a concrete example of this phenomenon. One sees that when $k_{1},k_{2},k_{3}$ are not all zero there is an obvious counterexample coming from Theorem 6.1. For a generalized strongly admissible sequence $s(t)$, set $\alpha=[s(t)^{\infty}]$ and $\beta=[s^{*}(\frac{1}{t})^{\infty}]$. By Theorems 2.20 and 6.1, we have $\mathcal{L}(\alpha)=\mathcal{L}(\beta)$. By Remark 4.5 (7), the periodic parts of $s(t)^{\infty}$ and $s^{*}(\frac{1}{t})^{\infty}$ are reverse-ordered with respect to each other, and in general they do not coincide unless $k_{1}=k_{2}=k_{3}=0$. By Serret’s theorem, two irrationals are $GL(2,\mathbb{Z})$-equivalent if and only if their periodic parts coincide, so $\alpha$ and $\beta$ need not be $GL(2,\mathbb{Z})$-equivalent. Moreover, there exist counterexamples that do not fall into this case. Such counterexample can be constructed using the counterexample of Question 7.3. To see this, set
$L=\frac{\sqrt{((3+0+1+2)81-2)^{2}-4}}{81}=\frac{2\sqrt{723}}{9}.$
Here, there exist at least two distinct quadratic irrationals $\alpha,\beta$ such that $L=\mathcal{L}(\alpha)=\mathcal{L}(\beta)$: one arising from the $(1,2,0)$-GM triple $(1,81,17)$, and another from the triple $(7,81,2)$. The triple $(1,81,17)$ corresponds to a vertex in $\mathrm{M}\mathbb{T}(1,2,0,\mathrm{id})$, where the fraction labeling of $81$ is $\tfrac{1}{3}$. The associated quadratic irrational is
$\alpha:=[(5,1,3,3,1,4)^{\infty}]=\frac{\sqrt{723}+25}{9}.$
On the other hand, the triple $(7,81,2)$ corresponds to a vertex in $\mathrm{M}\mathbb{T}(1,2,0,(1\ 2\ 3))$, where the fraction labeling of $81$ is $\tfrac{2}{3}$. The associated quadratic irrational is
$\beta:=[(5,1,1,5,3,2)^{\infty}]=\frac{\sqrt{723}+23}{9}.$
These periodic parts of their regular continued fraction expansions neither coincide nor are reversals of one another.
Each triple of a pair $(1,81,17)$ and $(7,81,2)$ that gives counterexample arises from distinct GM trees. This suggests that it is more appropriate to generalize the injectivity conjecture. The original injectivity conjecture is as follows:
Conjecture 7.5 ([aig]*Uniqueness conjecture IV).
Fix $\sigma\in\mathfrak{S}_{3}$ and consider $\mathrm{M}\mathbb{T}(0,0,0,\sigma)$. For any irreducible fractions $t,s\in[0,1]$, let $n_{t}$ and $n_{s}$ be Markov numbers with fraction labelings $t$ and $s$. If $n_{t}=n_{s}$, then $t=s$.
This conjecture is equivalent to Conjecture 7.1. The generalized version of this conjecture is as follows:
Conjecture 7.6.
Fix $\sigma\in\mathfrak{S}_{3}$ and consider $\mathrm{M}\mathbb{T}(k_{1},k_{2},k_{3},\sigma)$. For any irreducible fractions $t,s\in[0,1]$, let $n_{t}$ and $n_{s}$ be $(k_{1},k_{2},k_{3})$-GM numbers with fraction labelings $t$ and $s$. If $n_{t}=n_{s}$, then $t=s$.
When $k_{1}=k_{2}=k_{3}$, the tree $\mathrm{M}\mathbb{T}(k_{1},k_{2},k_{3},\sigma)$ is essentially independent of the permutation $\sigma$. In this case, the conjecture is equivalent to the statement that Question 7.3 for $k_{1}=k_{2}=k_{3}$ has a positive answer. By contrast, when $k_{1},k_{2},k_{3}$ are distinct, the tree $\mathrm{M}\mathbb{T}(k_{1},k_{2},k_{3},\sigma)$ does depend on $\sigma$, and the conjecture is no longer equivalent to Question 7.3. At present, no counterexamples to Conjecture 7.6 are known.
Regarding this conjecture, there exists a weaker version known as (generalized) Aigner conjecture, which has already been resolved when $k_{1}=k_{2}=k_{3}$:
Theorem 7.7 ([mcshane, llrs, banaian, banaian-sen]).
Fix $\sigma\in\mathfrak{S}_{3}$ and $k\in\mathbb{Z}_{\geq 0}$, and consider $\mathrm{M}\mathbb{T}(k,k,k,\sigma)$. For any irreducible fractions $t,s\in[0,1]$, let $n_{t}$ and $n_{s}$ be Markov numbers with fraction labelings $t$ and $s$. Then, we have the following:
  • (1)
    if $t$ and $s$ have the same numerator and $t<s$, then we have $n_{t}>n_{s}$,
  • (2)
    if $t$ and $s$ have the same denominator and $t<s$, then we have $n_{t}<n_{s}$,
  • (3)
    if $t$ and $s$ have the same sum of numerator and denominator and $t<s$, we have $n_{t}>n_{s}$.
However, when $k_{1},k_{2},k_{3}$ are not all equal, a counterexample can be found. For example, when $(k_{1},k_{2},k_{3},\sigma)$ = (0,0,1,(1 3 2))(0,0,1,(1\ 3\ 2)), the Table 6 shows that $n_{\frac{1}{4}}=34$ and $n_{\frac{2}{3}}=39$, which indicates that Theorem 7.7 (3) is not satisfied.